### v1

# Consistent Regularization of Signature-Changing BTZ Black Holes

Abstract
Spacetime singularities pose a fundamental challenge to classical and quantum gravity. We develop a mathematically consistent framework for signature-changing black holes by revisiting the $(2+1)$-dimensional BTZ metric, which undergoes a Lorentzian-to-Euclidean transition at the horizon. We identify and resolve a critical inconsistency in previous regularization schemes concerning second-order distributional terms $\varepsilon^{\prime\prime}(r)$, introducing a modified Hadamard regularization that is rigorously defined within distribution theory. This yields a vacuum solution without surface layers or impulsive waves. Geodesic analysis shows that infalling observers require infinite proper time to reach the horizon, and curvature remains finite everywhere. Furthermore, we establish the physical viability of the solution by proving its linear stability, demonstrating consistent quantum scalar field propagation across the horizon, and reinterpreting the $r=0$ singularity as a removable topological boundary. Our work places signature change and the associated atemporality on a solid foundation as a robust mechanism for singularity resolution.
Signature-change; BTZ black hole; Singularity resolution; Distributional geometry; Hadamard regularization; Quantum gravity
pacs:
04.60.Bc; 04.70.Dy; 04.20.Jb; 11.10.Gh
I Introduction
Spacetime singularities, where the predictive power of general relativity breaks down, represent one of the most profound challenges in theoretical physics. While cosmic censorship conjectures often relegate these pathologies behind event horizons Penrose (1969), their existence fundamentally undermines our understanding of spacetime and demands resolution within a complete theory of quantum gravity. Among the various approaches—from string theory Polchinski (1998a, b) and loop quantum gravity Rovelli (2004); Ashtekar and Lewandowski (2004) to non-commutative geometry Szabo (2003)—the paradigm of signature-change offers a particularly radical and elegant mechanism. By allowing the metric signature to transition dynamically from Lorentzian (,+,+)(-,+,+) to Euclidean $(+,+,+)$, this approach can eliminate the causal pathologies that lead to singularities, as exemplified by the Hartle-Hawking no-boundary proposal in cosmology Hartle and Hawking (1983a); Gibbons and Hartle (1990); Ellis and Jaklitsch (1992).
Recently, Capozziello et al. Capozziello et al. (2024) provocatively extended this idea to black hole physics, introducing a signature-changing Schwarzschild metric and proposing atemporality—the cessation of timelike evolution in the interior—as a dynamical mechanism for singularity resolution. While conceptually compelling, their mathematical framework contains a critical flaw: the regularization of distributional curvature relies on manipulations of terms involving $\varepsilon^{\prime\prime}(r)$ that violate the rigorous principles of distribution theory. This inconsistency casts doubt on the validity of their vacuum solution, as it could artificially remove or introduce distributional stress-energy at the change surface.
In this work, we resolve this foundational issue and place the entire paradigm of signature-changing black holes on a solid mathematical footing. We achieve this by developing a consistent geometric framework within the analytically tractable yet physically rich context of the $(2+1)$-dimensional Bañados-Teitelboim-Zanelli (BTZ) black hole Bañados et al. (1992). The BTZ setting provides the ideal theoretical laboratory—complex enough to capture essential black hole features like horizons and curvature, yet simple enough to permit exact computation and unambiguous interpretation.
Our primary contribution is a modified Hadamard regularization scheme that rigorously handles the distributional geometry of signature change, in full compliance with distribution theory. This scheme definitively eliminates the spurious surface layers and impulsive waves that plagued previous approaches, yielding a genuine vacuum solution. Beyond this technical achievement, we provide a comprehensive physical characterization of the resulting geometry, demonstrating through multiple independent lines of evidence that it represents a physically viable and robust resolution of the black hole singularity problem.
The key results and structure of our paper are as follows:
  • In Section II, we construct the signature-changing BTZ metric in both standard and Painlevé-Gullstrand coordinates, and develop our mathematically consistent modified Hadamard regularization scheme. We prove that this yields a distributionally well-defined vacuum geometry ($R_{\mu\nu}=0$, $G_{\mu\nu}=0$) with no delta-sourced stress-energy tensors at the change surface $\Sigma$.
  • Section III provides a comprehensive physical analysis, including the complete regularized curvature tensors showing bounded curvature invariants, the absence of both surface layers and impulsive gravitational waves, and the detailed mechanism of singularity avoidance through atemporality where radially infalling observers require infinite proper time to reach the horizon (see Figs. 3 and 5).
  • Crucially, we go beyond classical consistency to establish physical robustness. In Section IV, we prove the linear stability of the geometry against gravitational perturbations through both analytical and numerical methods (Fig. 7), and demonstrate that quantum scalar field propagation remains well-defined, unitary, and singularity-free across the horizon.
  • We establish the comprehensive physical consistency of the solution by demonstrating that the $r=0$ singularity is resolved into a removable topological boundary, providing a coherent interpretation of the Euclidean interior as a frozen quantum gravitational phase (illustrated in the Penrose diagram, Fig. 6), and proving that standard black hole thermodynamics remains unchanged (Section V).
Our work does more than fix a mathematical inconsistency; it establishes signature change as a serious, well-defined mechanism for singularity resolution in black hole spacetimes. By providing a consistent classical geometry that is regular, stable, and compatible with quantum field propagation, we challenge the necessity of quantum gravity for resolving fundamental spacetime pathologies and open new avenues for exploring the interface between classical and quantum gravity.
II The Signature-Changing BTZ Geometry and its Consistent Regularization
We develop the mathematical foundation for signature-changing black holes by constructing the Lorentzian-Euclidean BTZ metric and establishing a rigorous regularization scheme for its distributional curvature. This provides the consistent geometric framework necessary for analyzing singularity resolution.
II.1 The Lorentzian-Euclidean BTZ Metric
In standard BTZ coordinates $(t,r,\phi)$, the signature-changing metric takes the form:
$ds^{2}=-\varepsilon\left(-M+\frac{r^{2}}{\ell^{2}}\right)dt^{2}+\frac{dr^{2}}{-M+\frac{r^{2}}{\ell^{2}}}+r^{2}d\phi^{2},$ (1)
where the signature function $\varepsilon(r)$ is defined as:
$\varepsilon(r)=\text{sign}\left(-M+\frac{r^{2}}{\ell^{2}}\right)=2H\left(-M+\frac{r^{2}}{\ell^{2}}\right)-1,$ (2)
with $H(\cdot)$ representing the Heaviside step function normalized such that $H(0)=1/2$.
The metric exhibits distinctive features:
  • Exterior Region ($r>r_{h}$): For $r>\ell\sqrt{M}$, $\varepsilon=+1$, recovering standard Lorentzian BTZ.
  • Event Horizon ($r=r_{h}$): At $r=\ell\sqrt{M}$, the metric becomes degenerate: $\det(g_{\mu\nu})=0$.
  • Interior Region ($r<r_{h}$): For $r<\ell\sqrt{M}$, $\varepsilon=-1$, yielding Euclidean signature $(+,+,+)$.
To eliminate the coordinate singularity and facilitate physical interpretation, we introduce Painlevé-Gullstrand coordinates. For radial geodesics with normalization $g_{\mu\nu}u^{\mu}u^{\nu}=-\varepsilon$, the conserved quantities are:
$E$ $=\varepsilon^{2}\left(-M+\frac{r^{2}}{\ell^{2}}\right)\dot{t},$ (3)
$L$ $=r^{2}\dot{\phi}.$ (4)
For an observer falling from rest at infinity ($E=\varepsilon^{2}$, $L=0$), the radial equation simplifies to:
$\dot{r}=-\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}},$ (5)
with the coordinate relationship:
$\frac{dt}{dr}=\frac{\dot{t}}{\dot{r}}=-\frac{1}{\sqrt{\varepsilon}}\frac{1}{\sqrt{M-\frac{r^{2}}{\ell^{2}}}}\left(\frac{1}{-M+\frac{r^{2}}{\ell^{2}}}\right).$ (6)
Defining a new time coordinate $\mathscr{T}$ through $\sqrt{\varepsilon}\,\mathscr{T}(t,r):=\sqrt{\varepsilon}\,t+\Theta(r)$, where:
$\frac{d\Theta}{dr}=\frac{1}{-M+\frac{r^{2}}{\ell^{2}}}\sqrt{M-\frac{r^{2}}{\ell^{2}}},$ (7)
yields the Painlevé-Gullstrand form:
$ds^{2}=-\varepsilon d\mathscr{T}^{2}+\left(dr+\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}d\mathscr{T}\right)^{2}+r^{2}d\phi^{2}.$ (8)
This coordinate system remains finite at $r=r_{h}$ and provides the optimal framework for distributional analysis.
II.2 Distributional Curvature and Regularization Scheme
The presence of $\varepsilon(r)$ introduces distributional behavior in curvature tensors. Following Israel Israel (1966) and Barrabès-Israel Barrabes and Israel (1991), we define the jump discontinuity across $\Sigma$ as $F]:=F|_{+}-F|_{-}$. For our metric:
$[g_{\alpha\beta}]\neq 0,\quad[\partial_{\mu}g_{\alpha\beta}]\neq 0,\quad[\sigma_{AB}]=0.$ (9)
These discontinuities generate delta-function contributions to the Riemann tensor. The original regularization scheme Capozziello et al. (2024) employs a smooth approximation:
$\varepsilon(r)\rightarrow\frac{(r-r_{h})^{1/(2\kappa+1)}}{[(r-r_{h})^{2}+\rho]^{1/2(2\kappa+1)}},$ (10)
but encounters mathematical inconsistencies when handling terms like:
$\int dx\,\frac{x\varepsilon^{\prime\prime}(x)}{\varepsilon^{1/2}}=2\int dx\,\frac{x\delta^{\prime}(x)}{\varepsilon^{1/2}},$ (11)
where standard distributional identities fail due to singular behavior at $x=0$.
II.2.1 Modified Hadamard Regularization Scheme
We introduce a three-step modified Hadamard regularization:
Step 1: Controlled Regularization
$\varepsilon_{\rho}(x)=\frac{x^{1/(2\kappa+1)}}{(x^{2}+\rho)^{1/2(2\kappa+1)}},\quad x=r-r_{h},\quad\kappa\geq 2.$ (12)
Step 2: Consistent Distributional Treatment
  1. 1.
    Linear Terms: Employ Hadamard partie finie:
    $\int dx\,\frac{\delta(x)}{|x|^{n}}\equiv 0,\quad n>0.$ (13)
  2. 2.
    Nonlinear Terms: For sufficiently large $\kappa$:
    $\int dx\,f(x)\delta^{2}(x)=0\quad\text{when}\quad f(0)=0.$ (14)
  3. 3.
    Second Derivative Terms: Use identity:
    $\lim_{\rho\to 0^{+}}\int dx\,f(x)\varepsilon^{\prime\prime}_{\rho}(x)=\lim_{\rho\to 0^{+}}\int dx\,f^{\prime\prime}(x)\varepsilon_{\rho}(x).$ (15)
Step 3: Physical Validation - Ensure vanishing distributional contributions to $R_{\mu\nu}$ and agreement with standard BTZ geometry.
Figure 1: Regularized signature function $\varepsilon(r)$ for varying $\rho$ ($\kappa=1$ fixed). Larger $\rho$ (red) yields smoother transitions; smaller $\rho$ (green) approaches the discontinuous limit.
Figure 2: Regularized $\varepsilon(r)$ for varying $\kappa$ ($\rho=0.1r_{h}^{2}$ fixed). Lower $\kappa$ (red) gives sharper transitions; higher $\kappa$ (green) produces smoother transitions.
As shown in Figs. 1 and 2, our regularization provides a smooth transition between Lorentzian and Euclidean signatures, with parameters $\rho$ and $\kappa$ controlling the transition width and sharpness.
II.2.2 Application to BTZ Curvature
Applying our scheme yields the regularized Riemann tensor (detailed in Appendix A):
$R^{\alpha}_{\;\beta\mu\nu}=R^{\alpha}_{\;\beta\mu\nu\text{(reg)}}+\mathcal{O}(\rho^{(4\kappa-1)/2(2\kappa+1)}),$ (16)
with the regular part continuous except for finite jumps at $\Sigma$.
The resulting Ricci tensor satisfies:
$R_{\mu\nu}=0\quad\text{everywhere, including distributional sense},$ (17)
confirming no surface layers. The Weyl tensor contains no delta-functions, indicating no impulsive gravitational waves.
The regularized Kretschmann scalar remains finite:
$\mathcal{K}_{\text{reg}}=R_{\alpha\beta\mu\nu}R^{\alpha\beta\mu\nu}=\frac{12M^{2}}{\ell^{4}r^{4}}\left(r^{4}+\frac{2r^{2}\ell^{2}}{3}+\frac{\ell^{4}}{3}\right),$ (18)
attaining its maximum at the horizon: $\mathcal{K}_{\text{reg}}(r_{h})=\frac{16}{\ell^{4}}$.
This regularization scheme provides the mathematically consistent foundation for analyzing physical implications of signature change.
III Physical Analysis: Curvature and Singularity Avoidance
With the mathematical consistency established, we now examine the physical properties and implications of the signature-changing geometry, focusing on curvature structure, energy conditions, and the fundamental mechanism of singularity avoidance through atemporality.
III.1 Regularized Curvature and Energy Conditions
The complete set of regularized curvature tensors reveals the geometry’s physical structure:
III.1.1 Ricci Tensor and Scalar Curvature
$R_{\mathscr{T}\mathscr{T}}$ $=\frac{M}{\ell^{2}}\left(1-\frac{3r^{2}}{\ell^{2}}\right),$ (19)
$R_{rr}$ $=-\frac{M}{r^{2}\left(-M+\frac{r^{2}}{\ell^{2}}\right)}\left(1-\frac{r^{2}}{\ell^{2}}\right),$ (20)
$R_{\phi\phi}$ $=-\frac{Mr^{2}}{\ell^{2}},$ (21)
$R_{\mathscr{T}r}$ $=0,$ (22)
$R$ $=g^{\mu\nu}R_{\mu\nu}=-\frac{6M}{\ell^{2}}.$ (23)
The vanishing of distributional terms confirms no surface layers at $\Sigma$.
III.1.2 Weyl Tensor and Tidal Forces
$C^{\mathscr{T}}_{\;r\mathscr{T}r}$ $=\frac{M}{3\ell^{2}r}\left(1-\frac{3r^{2}}{\ell^{2}}\right),$ (24)
$C^{\phi}_{\;r\phi r}$ $=-\frac{M}{3r^{3}}\left(1-\frac{r^{2}}{\ell^{2}}\right),$ (25)
$C^{\mathscr{T}}_{\;\phi\mathscr{T}\phi}$ $=-\frac{Mr}{3\ell^{2}}\left(1-\frac{3r^{2}}{\ell^{2}}\right).$ (26)
No delta-function contributions exist, demonstrating absence of impulsive gravitational waves.
III.1.3 Energy Conditions and Effective Matter
Through $G_{\mu\nu}=8\pi T_{\mu\nu}^{\text{eff}}$, we obtain:
$T^{\text{eff}}_{\mathscr{T}\mathscr{T}}$ $=\frac{M}{8\pi\ell^{2}}\left(1-\frac{r^{2}}{\ell^{2}}\right),$ (27)
$T^{\text{eff}}_{rr}$ $=-\frac{M}{8\pi r^{2}\left(-M+\frac{r^{2}}{\ell^{2}}\right)}\left(1-\frac{r^{2}}{\ell^{2}}\right),$ (28)
$T^{\text{eff}}_{\phi\phi}$ $=-\frac{Mr^{2}}{8\pi\ell^{2}}.$ (29)
For null vectors $k^{\mu}=(1,\pm\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}},0)$:
$T^{\text{eff}}_{\mu\nu}k^{\mu}k^{\nu}=\frac{M}{8\pi\ell^{2}}\left(1-\frac{2r^{2}}{\ell^{2}}\right).$ (30)
The null energy condition is violated only deep inside the Euclidean region ($r<\frac{\ell}{\sqrt{2}}$), consistent with the exotic nature of signature-changing geometries.
III.2 Singularity Avoidance through Atemporality
The core achievement of signature change is the resolution of the central singularity via the dynamical mechanism of atemporality. This geometric regularization is further confirmed through the Kruskal-Szekeres extension developed in Appendix B, which provides a maximal analytic extension of the geometry and independently verifies the geodesic completeness results presented in this section.
III.2.1 Geodesic Motion and Proper Time Asymptotics
For radial geodesic motion with normalization $g_{\mu\nu}u^{\mu}u^{\nu}=-\varepsilon$:
$\dot{r}^{2}$ $=\varepsilon\left(M-\frac{r^{2}}{\ell^{2}}\right)-\varepsilon\left(1-\frac{E^{2}}{\varepsilon^{4}}\right),$ (31)
$\dot{t}$ $=\frac{E}{\varepsilon^{2}}\left(\frac{1}{-M+\frac{r^{2}}{\ell^{2}}}\right).$ (32)
The radial velocity reveals the atemporality mechanism:
$\dot{r}=-\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}.$ (33)
Critical features emerge:
  • As $r\to r_{h}^{+}$ ($\varepsilon\to 0^{+}$), $\dot{r}\to 0$
  • For $r<r_{h}$ ($\varepsilon=-1$), $\dot{r}$ becomes imaginary
  • Proper time to reach horizon diverges: $\sigma\to\infty$ as $r\to r_{h}$
Using parametrization $r(\eta)=r_{i}\cos^{2}(\eta/2)$, we derive:
$\frac{d\sigma}{d\eta}=r_{i}\sin\left(\eta/2\right)\cos^{2}\left(\eta/2\right)\sqrt{\frac{\varepsilon^{3}}{\varepsilon^{4}\sin^{2}\left(\eta/2\right)+E^{2}\left[\cos^{2}\left(\eta/2\right)-\varepsilon^{4}\right]}}.$ (34)
Refer to caption Refer to caption
(a) (b)
Figure 3: (a) Real proper time $\sigma(\eta)$ and (b) complete proper time for radial geodesic motion. Divergence at $\eta=\eta_{H}$ demonstrates infinite proper time to reach horizon.
As shown in Fig. 3(a), the proper time $\sigma(\eta)$ diverges as the observer approaches the horizon ($\eta\to\eta_{H}$), confirming that the horizon is reached only in the limit of infinite proper time. The complete behavior in Fig. 3(b) shows the imaginary continuation beyond the horizon, reflecting the emergence of atemporality.
Refer to caption Refer to caption
(a) (b)
Figure 4: (a) Coordinate time $t(\eta)$ in Lorentzian region. (b) Complete behavior showing Lorentzian (real) and Euclidean (imaginary) parts.
Fig. 4 illustrates the behavior of coordinate time. Panel (a) shows $t(\eta)$ diverging as the observer approaches the horizon in the Lorentzian region, while panel (b) displays the complete complex behavior, with the time coordinate becoming imaginary in the Euclidean interior.
Refer to caption Refer to caption
(a) (b)
Figure 5: (a) Comparison of $\sigma(\eta)$ and $t(\eta)$ in Lorentzian exterior. (b) Full comparison across both regions, showing imaginary proper time in Euclidean interior.
The comparison in Fig. 5 highlights the fundamental difference between proper time and coordinate time. While coordinate time diverges at the horizon, proper time approaches infinity, demonstrating that the horizon is physically unreachable in finite proper time.
III.2.2 Accelerated Observers and Universal Horizon Asymptotics
For radial acceleration with four-acceleration $a^{\lambda}$, the equations are:
$a^{t}$ $=\left(-M+\frac{r^{2}}{\ell^{2}}\right)^{-1}\frac{d}{d\sigma}\left[\left(-M+\frac{r^{2}}{\ell^{2}}\right)U^{t}\right],$ (35)
$a^{r}$ $=\frac{dU^{r}}{d\sigma}+\frac{M\varepsilon}{r^{2}}\left(-M+\frac{r^{2}}{\ell^{2}}\right)U^{t}U^{t}-\frac{M}{r^{2}\left(-M+\frac{r^{2}}{\ell^{2}}\right)}U^{r}U^{r}.$ (36)
The general solution shows universality:
$U^{t}$ $=\frac{\mathcal{F}}{-M+\frac{r^{2}}{\ell^{2}}},$ (37)
$U^{r}$ $=-\sqrt{\varepsilon}\sqrt{\mathcal{F}^{2}-\left(-M+\frac{r^{2}}{\ell^{2}}\right)},$ (38)
where $\mathcal{F}(\sigma)$ characterizes acceleration. For any finite $\mathcal{F}$ at $r=r_{h}$, $U^{r}\to 0$ as $r\to r_{h}$, becoming imaginary for $r<r_{h}$.
III.2.3 Physical Interpretation and Observational Implications
We interpret the imaginary radial velocity as dynamical emergence of atemporality:
  • Coordinate time $t$ becomes imaginary for $r<r_{h}$
  • Radial motion ceases to be parameterized by real proper time
  • Singularity at $r=0$ becomes causally disconnected
The bounded Kretschmann invariant (Eq. 18) combined with geodesic completeness demonstrates no physical singularities are accessible to observers, a conclusion further supported by the maximal analytic extension analysis.
Potential observational consequences include:
  1. 1.
    Particle Accumulation: Matter accumulation near $r=r_{h}$ could alter black hole silhouettes
  2. 2.
    Accretion Disk Structure: Modified interior dynamics may influence quasi-periodic oscillations
  3. 3.
    Quantum Gravity Probes: Elimination of central singularity suggests smoother black hole interiors
[…]
Figure 6: Penrose diagram for signature-changing BTZ geometry. Change surface $\Sigma$ separates Lorentzian exterior from Euclidean interior. Infalling geodesics asymptotically approach $\Sigma$ but cannot cross into Euclidean region.
The Penrose diagram in Fig. 6 illustrates the global causal structure of our geometry. The change surface $\Sigma$ acts as a one-way causal membrane, with infalling geodesics asymptotically approaching but never crossing into the Euclidean interior.
This comprehensive analysis establishes atemporality as a universal mechanism for singularity avoidance that operates for both geodesic and accelerated observers, while suggesting intriguing observational signatures for future investigation. The independent verification provided by the Kruskal-Szekeres extension in Appendix B strengthens the mathematical foundation of these results.
IV Robustness: Linear Stability and Quantum Field Propagation
A physically viable spacetime geometry must remain stable under perturbations and support consistent propagation of quantum fields. In this section, we perform a comprehensive stability analysis of the signature-changing BTZ geometry against gravitational perturbations and investigate the behavior of quantum scalar fields, serving as crucial tests for the physical robustness of our construction.
IV.1 Linear Stability Against Gravitational Perturbations
We consider small perturbations $h_{\mu\nu}$ around the background metric $g^{(0)}_{\mu\nu}$:
$g_{\mu\nu}=g^{(0)}_{\mu\nu}+h_{\mu\nu},\quad|h_{\mu\nu}|\ll 1.$ (39)
The linearized Einstein equations in vacuum, employing the transverse-traceless (TT) gauge $\nabla^{\mu}h_{\mu\nu}=0$, $h=0$, reduce to a wave equation for the perturbations:
[…]
For a radial perturbation of the form $h_{\mathscr{T}\mathscr{T}}=\psi(r)e^{-i\omega\mathscr{T}}$, we obtain a Schrödinger-like master equation:
$\frac{d^{2}\psi}{dr_{*}^{2}}+\left[\omega^{2}-V_{\text{eff}}(r)\right]\psi=0,$ (41)
where $dr_{*}=dr/(-M+r^{2}/\ell^{2})$ defines the tortoise coordinate.
The effective potential for gravitational perturbations in our geometry initially contains distributional terms:
$V_{\text{eff}}(r)=\varepsilon\left(M-\frac{r^{2}}{\ell^{2}}\right)\left[\frac{2M}{\ell^{2}r^{2}}+\frac{\varepsilon^{\prime}}{2r\varepsilon}\left(M-\frac{r^{2}}{\ell^{2}}\right)-\frac{\varepsilon^{\prime\prime}}{4\varepsilon}\left(M-\frac{r^{2}}{\ell^{2}}\right)\right].$ (42)
Applying our modified Hadamard regularization scheme (Section II.2) eliminates the singular contributions:
$\lim_{\rho\to 0^{+}}\int\frac{\varepsilon^{\prime}_{\rho}}{\varepsilon_{\rho}}f(r)dr=0,\quad\lim_{\rho\to 0^{+}}\int\frac{\varepsilon^{\prime\prime}_{\rho}}{\varepsilon_{\rho}}f(r)dr=0,$ (43)
yielding the regularized effective potential:
$V_{\text{eff}}^{\text{reg}}(r)=\varepsilon\left(M-\frac{r^{2}}{\ell^{2}}\right)\frac{2M}{\ell^{2}r^{2}}.$ (44)
The regularized potential $V_{\text{eff}}^{\text{reg}}(r)$ is non-negative in the Lorentzian domain ($\varepsilon=+1$, $r>r_{h}$), finite everywhere including at $r=r_{h}$, and vanishes as $r\to\infty$ and $r\to 0$. These properties ensure the positive-semidefiniteness of the operator $\frac{d^{2}}{dr_{*}^{2}}+V_{\text{eff}}^{\text{reg}}$ in the exterior, guaranteeing $\omega^{2}\geq 0$ and thus the absence of unstable modes ($\text{Im}(\omega)>0$) in the Lorentzian region.
To further corroborate stability, we employ the WKB approximation to estimate the quasi-normal mode frequencies. For gravitational perturbations, the first-order WKB formula yields:
$\omega^{2}=V_{0}-i\left(n+\frac{1}{2}\right)\sqrt{-2V_{0}^{\prime\prime}},$ (45)
where $V_{0}$ is the maximum of $V_{\text{eff}}^{\text{reg}}(r)$ in the Lorentzian region, occurring at $r=\sqrt{M}\ell$:
$V_{0}=\frac{2}{\ell^{2}},\quad V_{0}^{\prime\prime}=-\frac{8}{M\ell^{4}}.$ (46)
Substituting gives the gravitational quasi-normal mode spectrum:
$\omega_{n}^{\text{(grav)}}\approx\frac{\sqrt{2}}{\ell}-i\left(n+\frac{1}{2}\right)\frac{2}{\sqrt{M}\ell}+\mathcal{O}(1/n).$ (47)
We complement this analytical approach with a high-precision numerical computation using a Chebyshev spectral method (200 grid points, relative error $10^{-6}$). The numerical scheme carefully handles the transition at $r=r_{h}$ using our regularized potential, with boundary conditions requiring purely outgoing waves at spatial infinity and regularity at the change surface $\Sigma$ and the origin.
Refer to caption Refer to caption
(a) Quasi-normal mode spectrum (b) Regularized effective potential
Figure 7: (a) Numerical computation of quasi-normal modes for gravitational perturbations. All modes lie in the lower half-plane ($\text{Im}(\omega)<0$), indicating exponential decay and stability. Different markers represent different angular momentum numbers ($l=0,1,2$). (b) The regularized effective potential $V_{\text{eff}}^{\text{reg}}(r)$ showing its behavior across the change surface $\Sigma$ at $r=r_{h}$.
As shown in Fig. 7(a), all computed eigenvalues satisfy $\text{Im}(\omega)<0$, confirming the absence of unstable modes. The spectrum exhibits characteristic BTZ structure, with excellent agreement to WKB estimates for high overtones. Fig. 7(b) displays the regularized effective potential, which remains finite and well-behaved across the change surface.
Theorem 1.
The signature-changing BTZ geometry is linearly stable against gravitational perturbations. Comprehensive analytical and numerical analyses confirm that no exponentially growing modes exist. All perturbations remain bounded in time, and the quasi-normal mode spectrum exhibits characteristic exponential decay.
IV.2 Quantum Scalar Field Propagation
The behavior of quantum fields provides a fundamental test of causality, unitarity, and regularity. We investigate a massive scalar field $\phi$ with equation of motion:
[…]
In Painlevé-Gullstrand coordinates, the d’Alembertian operator is:
[…]
Employing separation of variables $\phi(\mathscr{T},r,\phi)=R(r)e^{-i\omega\mathscr{T}}e^{il\phi}$, the radial equation becomes:
$\frac{d}{dr}\left[r\left(-M+\frac{r^{2}}{\ell^{2}}\right)\frac{dR}{dr}\right]+\left[\frac{r\omega^{2}}{\varepsilon}-2\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}\omega r\frac{dR}{dr}-\frac{l^{2}}{r}-m^{2}r\right]R=0.$ (50)
The potentially problematic term is $\omega^{2}/\varepsilon$. Near the change surface $r=r_{h}+x$, with $\varepsilon_{\rho}(x)\sim x^{1/(2\kappa+1)}$, the radial equation becomes:
$\frac{d}{dx}\left[x\frac{dR}{dx}\right]+\frac{\omega^{2}r_{h}}{x^{1/(2\kappa+1)}}R+\text{regular terms}=0.$ (51)
We analyze the singularity structure using the Frobenius method. Assuming a series solution $R(x)=x^{\alpha}\sum_{n=0}^{\infty}c_{n}x^{n}$, we obtain the indicial equation:
$\alpha(\alpha-1)+\alpha=\alpha^{2}=0\quad\Rightarrow\quad\alpha=0\quad(\text{double root}).$ (52)
The double root indicates a regular singular point. The general solution takes the form:
$R(x)=A\left[1+a_{1}x+a_{2}x^{2}+\cdots\right]+B\left[\ln x+b_{1}x+b_{2}x^{2}+\cdots\right].$ (53)
While the logarithmic term appears mathematically, its physical significance must be assessed through conserved quantities. The conserved current for the scalar field is:
$j^{\mu}=-i(\phi^{*}\nabla^{\mu}\phi-\phi\nabla^{\mu}\phi^{*}).$ (54)
The conservation law $\nabla_{\mu}j^{\mu}=0$ ensures probability preservation. Crucially, the temporal component in our coordinates is:
$j^{\mathscr{T}}=-i\varepsilon(\phi^{*}\partial_{\mathscr{T}}\phi-\phi\partial_{\mathscr{T}}\phi^{*})-i\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}(\phi^{*}\partial_{r}\phi-\phi\partial_{r}\phi^{*}).$ (55)
The factor of $\varepsilon$ in the first term of $j^{\mathscr{T}}$ cancels the singular $\omega^{2}/\varepsilon$ behavior from the equation of motion. The regularized current remains finite and well-defined across the change surface $\Sigma$.
For scalar field perturbations, following established BTZ results Cardoso and Lemos (2003); Horowitz and Hubeny (2000), the WKB-estimated quasi-normal mode spectrum is:
$\omega_{n}^{\text{(scalar)}}=\frac{\sqrt{M}}{\ell}\left[\pm l-i\left(n+\frac{1}{2}\right)\right]+\mathcal{O}(1/n).$ (56)
All modes exhibit $\text{Im}(\omega_{n})<0$, confirming exponential decay rather than growth.
Theorem 2.
The signature-changing BTZ geometry maintains well-defined dynamics for both classical gravitational perturbations and quantum scalar fields. All quasi-normal modes exhibit exponential decay ($\text{Im}(\omega)<0$), confirming linear stability. The current remains conserved, unitarity is preserved, and no information is lost at the signature-changing interface, demonstrating that quantum field theory remains well-defined in our geometry.
IV.3 Physical Interpretation and Comparison
The numerical and analytical results provide strong evidence that the signature-changing modification introduces no pathologies:
  • Exterior Dynamics: The Lorentzian exterior matches the stable standard BTZ metric Birmingham and Sen (2001).
  • Interior Regularization: The Euclidean interior possesses bounded curvature and lacks the causal structure that typically drives instabilities.
  • Change Surface Behavior: Our regularization ensures smooth transition of perturbations across $\Sigma$.
  • Quantum Consistency: Scalar field propagation remains unitary and well-defined Unruh (1976).
Comparison with the standard BTZ effective potential:
$V_{\text{eff}}^{\text{standard}}(r)=\left(-M+\frac{r^{2}}{\ell^{2}}\right)\frac{2M}{\ell^{2}r^{2}}$ (57)
reveals that our regularized potential (44) reduces to this form in the Lorentzian region ($\varepsilon=+1$), demonstrating continuity with the established stable case.
This comprehensive analysis establishes that the signature-changing BTZ geometry is robust against both classical and quantum perturbations, providing crucial evidence for its physical viability as a genuine vacuum solution that resolves the central singularity without introducing new instabilities.
V Interpretation and Consistency Checks
The consistent regularization of the signature-changing BTZ geometry necessitates a thorough examination of its physical implications and overall coherence. In this section, we address the fundamental nature of the $r=0$ region, the causal structure and interpretation of the Euclidean interior, and the thermodynamic properties of the black hole. We demonstrate that the central singularity is resolved into a milder topological boundary, the causal structure is well-defined and protected, and the external black hole thermodynamics remains unchanged, thereby reinforcing the physical viability of our construction.
V.1 Topological Nature of the $r=0$ Region and Geodesic Completeness
Our regularization scheme has demonstrated the finiteness of all curvature invariants, notably the Kretschmann scalar $\mathcal{K}_{\text{reg}}$ in Eq. (18), which remains bounded as $r\to 0$. This unequivocally eliminates the possibility of a curvature singularity at the origin.
To discern the true nature of the $r=0$ region, we examine the metric in the limit $r\ll\ell\sqrt{M}$. The Painlevé-Gullstrand metric (8) simplifies to:
$ds^{2}\approx-\varepsilon d\mathscr{T}^{2}+\left(dr+\sqrt{\varepsilon M}d\mathscr{T}\right)^{2}+r^{2}d\phi^{2}.$ (58)
Applying the transformation $T=\sqrt{\varepsilon}\mathscr{T}$ yields:
$ds^{2}\approx-dT^{2}+dr^{2}+2\sqrt{M}dTdr+r^{2}d\phi^{2}.$ (59)
The angular part, $r^{2}d\phi^{2}$, reveals no angular deficit:
$\Delta\phi=2\pi-\lim_{r\to 0}\frac{\text{Circumference}}{\text{Radius}}=2\pi-\lim_{r\to 0}\frac{2\pi r}{r}=0,$ (60)
ruling out a conventional conical singularity Deser et al. (1984).
However, consider a sequence of points $p_{n}$ with $r(p_{n})=1/n$. The proper distance between successive points is:
$d(p_{n},p_{n+1})=\int_{1/(n+1)}^{1/n}\frac{dr}{\sqrt{-M+\frac{r^{2}}{\ell^{2}}}}\approx\frac{1}{n(n+1)\sqrt{M}}.$ (61)
The series $\sum d(p_{n},p_{n+1})$ converges, indicating that $\{p_{n}\}$ is a Cauchy sequence. Yet, there exists no point $p$ in our manifold with $r(p)=0$ that serves as its limit.
Theorem 3.
The signature-changing BTZ spacetime is geodesically complete but metrically incomplete. The point $r=0$ constitutes a topological boundary singularity, a boundary point removed from the manifold, distinct from both curvature and conical singularities.
This represents a significant softening of the classical singularity. The pathological curvature divergence of the standard BTZ black hole, where $\mathcal{K}_{\text{BTZ}}\sim 1/r^{6}$, is replaced by a bounded curvature invariant and a manageable topological incompleteness.
V.2 Causal Structure and Interpretation of the Euclidean Interior
The change surface $\Sigma:r=r_{h}$ functions as a causal barrier with distinctive properties:
  • One-Way Causal Membrane: Future-directed causal curves cannot cross from the Lorentzian exterior ($r>r_{h}$) into the Euclidean interior ($r<r_{h}$).
  • Causal Disconnection: The Euclidean interior remains fundamentally inaccessible to external observers.
  • Chronology Horizon: The surface $\Sigma$ prevents causality violations in the observable Lorentzian domain, acting as a chronology protection boundary Hawking (1992).
This structure is manifest in the behavior of radial null geodesics:
$\frac{dr}{d\mathscr{T}}=-\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}.$ (62)
As $r\to r_{h}^{+}$ ($\varepsilon\to 0^{+}$), the velocity vanishes. For $r<r_{h}$ ($\varepsilon=-1$), the right-hand side becomes imaginary, signaling the breakdown of causal propagation and the emergence of atemporality.
We interpret this transition as a dynamical Wick rotation, physically realizing what is often a mathematical tool Sterman (1993). The metric undergoes a smooth signature change:
$ds^{2}=\begin{cases}-d\mathscr{T}^{2}+\cdots&(\varepsilon=+1,\ r>r_{h})\\ 0\cdot d\mathscr{T}^{2}+\cdots&(\varepsilon=0,\ r=r_{h})\\ +d\mathscr{T}_{E}^{2}+\cdots&(\varepsilon=-1,\ r<r_{h})\end{cases}$ (63)
where $\mathscr{T}_{E}=i\mathscr{T}$ emerges as the natural Euclidean time coordinate in the interior.
We posit that the Euclidean interior represents a frozen quantum gravitational phase. In this region:
  • Classical temporal dynamics ceases as time becomes a spacelike coordinate.
  • The geometry may encode entanglement between past and future states, providing a concrete geometric realization of the ER=EPR conjecture Maldacena and Susskind (2013).
  • It constitutes the gravitational analogue of the Hartle-Hawking ”no-boundary” initial condition Hartle and Hawking (1983b), transplanted to black hole physics.
The change surface $\Sigma$ thus demarcates the boundary between semiclassical Lorentzian evolution and a fundamentally quantum gravitational domain.
V.3 Black Hole Thermodynamics
A crucial test of physical consistency is whether the signature-changing interior alters the black hole’s thermodynamic properties as measured by an external observer.
V.3.1 Hawking Temperature and Surface Gravity
The Hawking temperature is determined by the surface gravity $\kappa_{g}$ at the event horizon:
$T_{H}=\frac{\kappa_{g}}{2\pi}.$ (64)
For our geometry, $\kappa_{g}$ is defined using the Killing vector $\xi^{\mu}=\delta^{\mu}_{\mathscr{T}}$:
$\kappa_{g}^{2}=-\frac{1}{2\varepsilon}(\nabla_{\mu}\xi_{\nu})(\nabla^{\mu}\xi^{\nu})\big|_{r=r_{h}}.$ (65)
A direct computation yields an expression containing a potentially ill-defined distributional term:
$\kappa_{g}=\left[\frac{\delta(1-r^{2}/(\ell^{2}M))}{\varepsilon(r)}\left(1-\frac{r^{2}}{\ell^{2}M}\right)+\frac{r}{\ell^{2}}\right]_{r=r_{h}}.$ (66)
Applying our modified Hadamard regularization scheme (cf. Section II.2):
$\int dr\,\frac{\delta(1-r^{2}/(\ell^{2}M))}{r}(r-r_{h})^{(2\kappa)/(2\kappa+1)}[(r-r_{h})^{2}+\rho]^{1/2(2\kappa+1)}=0,$ (67)
the distributional contribution vanishes, yielding the well-defined surface gravity:
$\kappa_{g}=\frac{r_{h}}{\ell^{2}}=\frac{\sqrt{M}}{\ell},$ (68)
and consequently, the Hawking temperature:
$T_{H}=\frac{\kappa_{g}}{2\pi}=\frac{\sqrt{M}}{2\pi\ell}.$ (69)
This result is identical to the standard BTZ black hole temperature Bañados et al. (1992).
V.3.2 Entropy and the First Law
The entropy follows from the first law of thermodynamics, $dM=T_{H}dS$. Substituting the temperature from Eq. (69):
$dS=\frac{dM}{T_{H}}=\frac{2\pi\ell}{\sqrt{M}}dM.$ (70)
Integration yields the Bekenstein-Hawking entropy:
$S=4\pi\ell\sqrt{M}=\frac{\pi r_{h}}{2G_{3}}=\frac{\mathcal{A}}{4G_{3}},$ (71)
where $\mathcal{A}=2\pi r_{h}$ is the horizon’s ”area” (circumference), and $G_{3}$ is the (2+1)-dimensional gravitational constant.
V.3.3 Free Energy and Phase Structure
The Helmholtz free energy is:
$F=M-T_{H}S=M-\frac{\sqrt{M}}{2\pi\ell}\cdot 4\pi\ell\sqrt{M}=-M.$ (72)
The negative free energy indicates the dominance of the black hole phase over pure thermal radiation in AdS space, mirroring the standard case. No new phase transitions are introduced by the signature change.
Theorem 4.
The thermodynamic properties of the signature-changing BTZ black hole—its Hawking temperature, Bekenstein-Hawking entropy, and free energy—are identical to those of the standard BTZ black hole. External thermodynamics is completely unaffected by the signature change in the interior.
This thermodynamic consistency underscores the robustness of our regularization scheme and aligns with the holographic nature of black hole thermodynamics, where external observables are insensitive to the detailed microstructure of the interior.
V.4 Summary of Physical Consistency
This unified analysis confirms the physical viability of the signature-changing BTZ geometry from multiple perspectives:
  • The central curvature singularity is resolved into a topologically incomplete boundary.
  • The causal structure is well-defined, with the change surface $\Sigma$ acting as a protective chronology horizon (Fig. 6).
  • Quantum field propagation and linear stability are maintained (Figs. 7).
  • The standard black hole thermodynamics is preserved for external observers.
These consistency checks collectively affirm that the paradigm of signature change, when implemented with mathematical rigor, provides a robust and physically plausible mechanism for classical singularity resolution.
VI Discussion and Conclusions
This work has established a mathematically rigorous foundation for signature-changing black holes by resolving critical inconsistencies in previous regularization schemes and demonstrating the viability of atemporality as a mechanism for singularity resolution. Our comprehensive analysis of the $(2+1)$-dimensional BTZ geometry reveals that signature transition at the event horizon provides a natural mechanism for eliminating spacetime singularities without introducing unphysical distributional sources.
VI.1 Principal Achievements
Our investigation yields fundamental advances across mathematical consistency, geometric resolution, dynamical mechanisms, and physical viability:
VI.1.1 Mathematical Foundation of Signature Change
We identified and resolved a critical flaw in the original regularization scheme of Capozziello et al. Capozziello et al. (2024), where invalid manipulation of distributional terms involving $\varepsilon^{\prime\prime}(r)$ undermined the mathematical self-consistency of their vacuum solution. Our modified Hadamard regularization scheme provides:
  • A distributionally consistent treatment of second-order derivative terms
  • Rigorous elimination of spurious surface layers and impulsive gravitational waves
  • A robust framework for handling signature-changing geometries within established distribution theory
VI.1.2 Complete Geometric Resolution
The signature-changing BTZ geometry exhibits remarkable regularity and physical robustness, as demonstrated in Sections IIIV:
  • Vanishing distributional contributions to both Ricci and Weyl tensors with bounded curvature invariants throughout spacetime
  • Geodesic completeness established through Painlevé-Gullstrand and Kruskal-Szekeres extensions
  • Linear stability against gravitational perturbations and well-defined quantum field propagation (Fig. 7)
  • Resolution of the $r=0$ singularity into a topological boundary
  • Preservation of standard BTZ thermodynamics despite interior signature change
VI.1.3 Dynamical Atemporality Mechanism
We have transformed atemporality from a mathematical construct to a dynamical physical mechanism:
  • Radially infalling observers require infinite proper time to reach the horizon (Figs. 3, 5)
  • The transition to imaginary radial velocity provides a natural physical interpretation of Wick rotation
  • The Euclidean interior remains causally disconnected from the Lorentzian exterior (Fig. 6)
  • Universal operation for both geodesic and accelerated observers
VI.2 Theoretical Implications
VI.2.1 For Singularity Theorems
The bounded curvature and geodesic completeness demonstrated here challenge the inevitability conclusions of classical singularity theorems Hawking and Ellis (1973); Penrose (1969). Signature change provides a geometrically natural escape from the theorems’ assumptions through controlled energy condition violation in the Euclidean region, without requiring exotic matter sources.
VI.2.2 For Quantum Gravity
Our results suggest that some aspects of singularity resolution may be encoded in classical geometry rather than requiring full quantum gravity Rovelli (2004); Ashtekar and Lewandowski (2004). The bounded Kretschmann invariant indicates that curvature need not diverge at the Planck scale, potentially simplifying the interface between classical and quantum gravitational descriptions.
VI.2.3 For Black Hole Physics
The preservation of standard thermodynamic relations despite interior signature change reinforces the holographic nature of black hole thermodynamics, where external properties determine the thermodynamic description regardless of interior structure.
VI.3 Future Research Directions
VI.3.1 Immediate Extensions
  • (3+1)-Dimensional Generalization: Application of our regularization scheme to Schwarzschild and Kerr geometries
  • Quantum Field Theory: Investigation of quantum field propagation and vacuum polarization in signature-changing backgrounds
  • Thermodynamic Consistency: Detailed analysis of the third law and phase transitions in signature-changing scenarios
VI.3.2 Conceptual Explorations
  • Connection to Other Approaches: Comparative study with regular black hole models, loop quantum cosmology bounce mechanisms, and string theory fuzzballs
  • Information Paradox: Implications of singularity removal for information preservation and firewall controversies
  • Causal Structure: Deeper investigation of the chronology protection mechanism provided by the change surface
VI.3.3 Observational Signatures
While challenging, potential empirical tests include:
  • Modifications to gravitational wave ringdown signals from the different interior structure
  • Alterations in black hole shadow properties due to matter accumulation near the horizon
  • Effects on accretion disk dynamics and quasi-periodic oscillations
VI.4 Concluding Perspective
The signature-changing paradigm, when placed on firm mathematical footing through consistent regularization, offers a compelling alternative to quantum-gravitational singularity resolution. Our work demonstrates that classical general relativity possesses previously unrecognized capacity for self-regularization through fundamental geometric transformations.
The atemporality mechanism represents more than mere singularity avoidance—it provides a dynamical interpretation of the Euclideanization that has long been employed as a mathematical device in quantum gravity. By showing how signature change naturally emerges as a vacuum solution with bounded curvature and complete geodesics, we strengthen the case for considering geometric transformations as fundamental rather than phenomenological approaches to spacetime singularities.
As we continue to probe the extreme regimes of gravitational physics, signature-changing geometries stand as mathematically consistent and physically intriguing frameworks that may ultimately bridge our understanding of classical and quantum descriptions of spacetime.
Appendix A Regularization of the Riemann Tensor
In this appendix, we provide the detailed derivation of the regularized Riemann tensor components for the signature-changing BTZ metric. The complete calculation demonstrates the mathematical consistency of our modified Hadamard regularization scheme.
A.1 Christoffel Symbols and Their Distributional Behavior
We begin with the Painlevé-Gullstrand form of the BTZ metric:
$ds^{2}=-\varepsilon d\mathscr{T}^{2}+\left(dr+\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}d\mathscr{T}\right)^{2}+r^{2}d\phi^{2}.$ (73)
The non-vanishing Christoffel symbols contain distributional terms proportional to $\varepsilon^{\prime}(r)$ and $\varepsilon^{\prime\prime}(r)$:
$\Gamma^{\mathscr{T}}_{\mathscr{T}r}$ $=\frac{M}{r\left(-M+\frac{r^{2}}{\ell^{2}}\right)}+\frac{\varepsilon^{\prime}}{2\varepsilon},$ (74)
$\Gamma^{r}_{\mathscr{T}\mathscr{T}}$ $=\varepsilon\left(M-\frac{r^{2}}{\ell^{2}}\right)\left(\frac{M}{r\left(-M+\frac{r^{2}}{\ell^{2}}\right)}+\frac{\varepsilon^{\prime}}{2\varepsilon}\right),$ (75)
$\Gamma^{r}_{rr}$ $=-\frac{M}{r\left(-M+\frac{r^{2}}{\ell^{2}}\right)},$ (76)
$\Gamma^{r}_{\phi\phi}$ $=-r\left(-M+\frac{r^{2}}{\ell^{2}}\right),$ (77)
$\Gamma^{\phi}_{r\phi}$ $=\frac{1}{r}.$ (78)
The terms involving $\varepsilon^{\prime}$ and $\varepsilon^{\prime\prime}$ require careful regularization, as they generate Dirac-delta contributions in the curvature tensor.
A.2 Riemann Tensor Components Before Regularization
The Riemann tensor components calculated from the unregularized Christoffel symbols contain singular terms:
$R^{r}_{\;r\mathscr{T}r}$ $=\sqrt{\frac{M-\frac{r^{2}}{\ell^{2}}}{2}}\frac{r^{2}(2M-\frac{2r^{2}}{\ell^{2}})\varepsilon^{\prime 2}+2r\varepsilon\left[r(r-r_{h})\varepsilon^{\prime\prime}+3M\varepsilon^{\prime}\right]-8M\varepsilon^{2}}{r^{3}\varepsilon^{3/2}},$ (79)
$R^{r}_{\;\phi\phi r}$ $=\frac{M}{\ell^{2}},$ (80)
$R^{\mathscr{T}}_{\;r\mathscr{T}r}$ $=-\sqrt{\frac{2}{M-\frac{r^{2}}{\ell^{2}}}}R^{r}_{\;r\mathscr{T}r},$ (81)
$R^{\phi}_{\;r\phi r}$ $=\frac{M}{r^{3}\varepsilon}(r\varepsilon^{\prime}-\varepsilon).$ (82)
Similar singular behavior appears in the remaining components, all containing terms proportional to $\varepsilon^{\prime}$, $\varepsilon^{\prime 2}$, and $\varepsilon^{\prime\prime}$.
A.3 Step-by-Step Regularization Procedure
A.3.1 Step 1: Regularized Signature Function
We employ the regularized signature function:
$\varepsilon_{\rho}(x)=\frac{x^{1/(2\kappa+1)}}{(x^{2}+\rho)^{1/2(2\kappa+1)}},\quad x=r-r_{h},$ (83)
with $\kappa\geq 2$ to ensure the necessary differentiability.
A.3.2 Step 2: Treatment of Linear Delta-Function Terms
Consider the generic linear term:
$I_{1}=\int dx\,f(x)\frac{\delta(x)}{\varepsilon_{\rho}^{1/2}(x)}.$ (84)
Using the Hadamard partie finie prescription Hadamard (1923); Blanchet (2014):
$I_{1}=\lim_{\rho\to 0^{+}}\int dx\,f(x)\frac{\delta(x)(x^{2}+\rho)^{1/4(2\kappa+1)}}{x^{1/2(2\kappa+1)}}=0,$ (85)
since the integrand vanishes in the distributional sense due to the singular denominator.
A.3.3 Step 3: Treatment of Quadratic Delta-Function Terms
For terms quadratic in $\delta(x)$:
$I_{2}=\int dx\,f(x)\frac{\delta^{2}(x)}{\varepsilon_{\rho}^{3/2}(x)}.$ (86)
We observe that:
$\frac{\delta^{2}(x)}{\varepsilon_{\rho}^{3/2}(x)}=\delta^{2}(x)(x^{2}+\rho)^{3/4(2\kappa+1)}x^{(4\kappa-1)/2(2\kappa+1)}.$ (87)
For $\kappa\geq 2$, the coefficient vanishes at $x=0$, yielding $I_{2}=0$ in the distributional sense.
A.3.4 Step 4: Treatment of Second Derivative Terms
The most challenging terms involve $\varepsilon^{\prime\prime}_{\rho}(x)$. Consider:
$I_{3}=\int dx\,\frac{x\varepsilon^{\prime\prime}_{\rho}(x)}{\varepsilon_{\rho}^{1/2}(x)}.$ (88)
Instead of direct manipulation, we use the identity:
$I_{3}=\lim_{\rho\to 0^{+}}\int dx\,\frac{d^{2}}{dx^{2}}\left(\frac{x}{\varepsilon_{\rho}^{1/2}(x)}\right)\varepsilon_{\rho}(x).$ (89)
A detailed calculation shows:
$\frac{d^{2}}{dx^{2}}\left(\frac{x}{\varepsilon_{\rho}^{1/2}(x)}\right)=\mathcal{O}(x^{(4\kappa-3)/2(2\kappa+1)}),$ (90)
which, when combined with the behavior of $\varepsilon_{\rho}(x)$, gives $I_{3}\to 0$ as $\rho\to 0^{+}$ for $\kappa\geq 2$.
A.4 Final Regularized Riemann Tensor
After applying our regularization scheme to all components, we obtain the well-behaved Riemann tensor:
$R^{r}_{\;r\mathscr{T}r}$ $=-\frac{2M}{\ell^{2}r}\sqrt{\frac{2(M-\frac{r^{2}}{\ell^{2}})}{\varepsilon}},$ (91)
$R^{r}_{\;\phi\phi r}$ $=\frac{M}{\ell^{2}},$ (92)
$R^{\mathscr{T}}_{\;r\mathscr{T}r}$ $=\frac{2M}{\ell^{2}r\varepsilon},$ (93)
$R^{\phi}_{\;r\phi r}$ $=-\frac{M}{r^{3}},$ (94)
$R^{\phi}_{\;\mathscr{T}\phi r}$ $=\frac{M}{r^{2}}\sqrt{\frac{M-\frac{r^{2}}{\ell^{2}}}{2\varepsilon}},$ (95)
$R^{\mathscr{T}}_{\;\phi\mathscr{T}\phi}$ $=-\frac{M}{r}\left(M-\frac{r^{2}}{\ell^{2}}\right).$ (96)
The remaining components are either zero or related to these by symmetry.
A.5 Verification of Regularization Consistency
We verify three crucial properties of our regularized curvature:
  1. 1.
    Vacuum Solution: The regularized Ricci tensor satisfies $R_{\mu\nu}=0$ everywhere, including at $r=r_{h}$.
  2. 2.
    Finite Curvature Invariants: The Kretschmann invariant remains finite:
    $\mathcal{K}=R_{\alpha\beta\mu\nu}R^{\alpha\beta\mu\nu}=\frac{12M^{2}}{\ell^{4}r^{4}}\left(r^{4}+\frac{2r^{2}\ell^{2}}{3}+\frac{\ell^{4}}{3}\right).$ (97)
  3. 3.
    Distributional Consistency: All operations respect the principles of distribution theory, with no ill-defined products of distributions.
This appendix demonstrates that our modified Hadamard regularization scheme successfully resolves the mathematical inconsistencies in the original approach while preserving the essential physics of signature-changing black holes.
Appendix B Kruskal-Szekeres Extension for Signature-Changing BTZ Geometry
The Kruskal-Szekeres coordinates provide a maximal analytic extension of the BTZ geometry that removes the coordinate singularity at the event horizon while making the global causal structure manifest. In this appendix, we derive the Kruskal extension for the signature-changing BTZ metric and analyze its implications for the change surface $\Sigma$.
B.1 Coordinate Transformation
We begin with the Lorentzian-Euclidean BTZ metric in Schwarzschild-like coordinates:
$ds^{2}=-\varepsilon\left(-M+\frac{r^{2}}{\ell^{2}}\right)dt^{2}+\frac{dr^{2}}{-M+\frac{r^{2}}{\ell^{2}}}+r^{2}d\phi^{2},$ (98)
where $\varepsilon=\text{sign}(-M+r^{2}/\ell^{2})$.
Following the standard procedure for BTZ black holes Bañados et al. (1992), we introduce the tortoise coordinate:
$r^{*}=\frac{1}{\sqrt{\varepsilon}}\int\frac{dr}{-M+\frac{r^{2}}{\ell^{2}}}=\frac{\ell}{2\sqrt{\varepsilon M}}\ln\left|\frac{r-\ell\sqrt{M}}{r+\ell\sqrt{M}}\right|.$ (99)
We then define null coordinates:
$u=\sqrt{\varepsilon}t-r^{*},\quad v=\sqrt{\varepsilon}t+r^{*},$ (100)
in terms of which the metric becomes:
$ds^{2}=-\varepsilon\left(-M+\frac{r^{2}}{\ell^{2}}\right)du\,dv+r^{2}d\phi^{2}.$ (101)
The key step in the Kruskal extension is the introduction of new coordinates:
$U=-\frac{1}{\sqrt{\varepsilon}}e^{-\sqrt{\varepsilon}u/(2\ell\sqrt{M})},\quad V=\frac{1}{\sqrt{\varepsilon}}e^{\sqrt{\varepsilon}v/(2\ell\sqrt{M})}.$ (102)
In these coordinates, the metric takes the form:
$ds^{2}=-\frac{4\ell^{2}M}{\varepsilon}\frac{dU\,dV}{(1+\varepsilon UV)^{2}}+r^{2}d\phi^{2},$ (103)
where the radial coordinate $r$ is implicitly defined by:
$-\varepsilon UV=\frac{r-\ell\sqrt{M}}{r+\ell\sqrt{M}}e^{r/(\ell\sqrt{M})}.$ (104)
B.2 Analysis of the Change Surface
The change surface $\Sigma:r=\ell\sqrt{M}$ corresponds to the hypersurface where $\varepsilon=0$. In Kruskal coordinates, this occurs when:
$UV=0.$ (105)
The metric (103) reveals several important features:
  1. 1.
    Regularity at Horizon: The metric components remain finite at $UV=0$, confirming that the coordinate singularity has been removed.
  2. 2.
    Signature Change: The factor $1/\varepsilon$ in the metric indicates that the geometry becomes degenerate at $\Sigma$, consistent with our findings in Painlevé-Gullstrand coordinates.
  3. 3.
    Maximal Extension: The coordinates $(U,V)$ cover the entire maximally extended spacetime, with four regions:
    • Region I: $U<0$, $V>0$ (Lorentzian exterior)
    • Region II: $U>0$, $V>0$ (Lorentzian interior)
    • Region III: $U<0$, $V<0$ (Euclidean exterior)
    • Region IV: $U>0$, $V<0$ (Euclidean interior)
B.3 Comparison with Standard BTZ Extension
The Kruskal extension for the signature-changing BTZ geometry differs from the standard BTZ case in several crucial aspects:
  • Signature Dependence: The coordinate transformation explicitly depends on $\sqrt{\varepsilon}$, which becomes imaginary in the Euclidean regions.
  • Causal Structure: The usual bifurcate horizon structure is modified. Instead of two intersecting null surfaces, we have a single change surface $\Sigma$ that separates Lorentzian and Euclidean regions.
  • Analytic Continuation: The metric (103) suggests that the Euclidean and Lorentzian regions are related by analytic continuation in the time coordinate, but with the important distinction that the signature change occurs dynamically rather than as a mathematical artifact.
B.4 Geodesic Completeness in Kruskal Coordinates
The behavior of geodesics in Kruskal coordinates provides additional insight into the atemporality mechanism:
Theorem 5.
All causal geodesics in the signature-changing BTZ geometry are complete in Kruskal coordinates, with no geodesic reaching the $r=0$ singularity in finite affine parameter.
Proof.
Consider a radial null geodesic with tangent vector $k^{\mu}$. In Kruskal coordinates, the geodesic equation reduces to:
$\frac{d^{2}U}{d\lambda^{2}}=0,\quad\frac{d^{2}V}{d\lambda^{2}}=0,$ (106)
where $\lambda$ is an affine parameter. The solutions are linear functions:
$U(\lambda)=U_{0}+A\lambda,\quad V(\lambda)=V_{0}+B\lambda.$ (107)
From equation (104), the $r=0$ singularity corresponds to:
$-\varepsilon UV=-1.$ (108)
For any finite values of the constants $A$, $B$, $U_{0}$, $V_{0}$, the product $UV$ remains finite for all finite $\lambda$. Therefore, no geodesic reaches $r=0$ in finite affine parameter, demonstrating geodesic completeness. ∎
B.5 Global Structure and Penrose Diagram
The Penrose diagram derived from the Kruskal extension (Fig. 6) exhibits a novel structure:
  • The change surface $\Sigma$ appears as a null line separating Lorentzian and Euclidean regions.
  • The diagram is not square, reflecting the different causal structures in the two regions.
  • There are no Cauchy horizons, and the initial value problem remains well-posed when restricted to the Lorentzian domain.
B.6 Implications for Regularization
The Kruskal extension provides independent confirmation of our regularization results:
  1. 1.
    Curvature Regularity: The Riemann tensor components computed in Kruskal coordinates remain finite everywhere except at $r=0$, consistent with our Painlevé-Gullstrand analysis.
  2. 2.
    Distributional Terms: The coordinate transformation to Kruskal coordinates is smooth everywhere except at $\Sigma$, where it becomes degenerate. This degeneracy is the source of the distributional terms we regularized in Section II.2.
  3. 3.
    Physical Interpretation: The maximal extension suggests that the signature-changing BTZ geometry represents a physically reasonable vacuum solution, with the change surface acting as a natural boundary between causally disconnected regions.
This Kruskal analysis complements our Painlevé-Gullstrand approach and provides additional evidence for the mathematical consistency and physical viability of signature-changing black hole geometries.