### v3

# Consistent Regularization of Signature-Changing BTZ Black Holes

Abstract
Spacetime singularities represent a fundamental challenge in gravitational physics. We present a mathematically consistent framework for signature-changing black holes based on the $(2+1)$-dimensional BTZ metric, where the signature transitions from Lorentzian (,+,+)(-,+,+) to Euclidean $(+,+,+)$ at the horizon. We identify and rectify a critical inconsistency in previous regularization schemes concerning second-order distributional terms $\varepsilon^{\prime\prime}(r)$, introducing a modified Hadamard regularization that respects distribution theory. This produces a vacuum solution free of surface layers and impulsive gravitational waves. Geodesic analysis reveals that radially infalling observers require infinite proper time to reach the horizon, effectively preventing access to the would-be singularity while maintaining finite curvature invariants throughout the spacetime. We further establish the physical robustness of the geometry by demonstrating linear stability against gravitational perturbations, showing that quantum scalar field propagation remains unitary and well-defined across the signature change, and reinterpreting the $r=0$ region as a topological boundary rather than a curvature singularity. Our work establishes atemporality via signature change as a mathematically rigorous mechanism for singularity avoidance in black hole spacetimes.
Signature-change; BTZ black hole; singularity regularization; Distributional geometry; Hadamard regularization; Quantum gravity
pacs:
04.60.Bc; 04.70.Dy; 04.20.Jb; 11.10.Gh
I Introduction
Spacetime singularities, where classical predictability breaks down, represent a central challenge in gravitational physics. While cosmic censorship conjectures typically hide these pathologies behind event horizons Penrose (1969), their existence demands resolution within a consistent framework for quantum gravity. Among various approaches—from string theory Polchinski (1998a, b) and loop quantum gravity Rovelli (2004); Ashtekar and Lewandowski (2004) to noncommutative geometry Szabo (2003)—the paradigm of signature change offers a geometric alternative: by allowing the metric to transition from Lorentzian (,+,+)(-,+,+) to Euclidean $(+,+,+)$, causal pathologies leading to singularities can be avoided, as exemplified in cosmology by the Hartle-Hawking no-boundary proposal Hartle and Hawking (1983a); Gibbons and Hartle (1990); Ellis and Jaklitsch (1992).
Recently, this idea was extended to black hole physics by Capozziello et al. Capozziello et al. (2024), who introduced a signature-changing Schwarzschild metric and proposed atemporality as a mechanism for singularity avoidance. However, their regularization scheme contains a critical mathematical inconsistency: terms involving $\varepsilon^{\prime\prime}(r)$ are manipulated in ways that violate distribution theory, potentially introducing or removing spurious stress-energy at the change surface. This undermines the claim of a vacuum solution and leaves open essential questions of stability, quantum field propagation, and thermodynamic consistency.
In this work, we resolve these issues and establish a mathematically rigorous foundation for signature-changing black holes. Our primary contributions are:
  1. 1.
    A modified Hadamard regularization scheme that respects distribution theory, properly handles $\varepsilon^{\prime\prime}(r)$ terms, and yields a genuine vacuum solution without surface layers or impulsive gravitational waves.
  2. 2.
    A complete physical analysis of the resulting $(2+1)$-dimensional BTZ geometry Bañados et al. (1992), demonstrating bounded curvature, geodesic completeness (infinite proper time to reach the horizon), linear stability, unitary quantum field propagation, and preservation of standard black hole thermodynamics.
  3. 3.
    A clarification of the avoidance mechanism: infinite proper time to cross the horizon constitutes singularity avoidance through atemporality, where the interior becomes Euclidean and causally disconnected from the exterior.
The BTZ setting provides an ideal theoretical laboratory—sufficiently rich to capture essential black hole features, yet tractable enough for exact computation and unambiguous interpretation. We show that signature change, when implemented with mathematical consistency, provides a robust classical mechanism for singularity avoidance, challenging the notion that quantum gravity is strictly necessary for this task.
The paper is structured as follows: Section II presents the signature-changing BTZ metric and our regularization scheme. Section III analyzes curvature and geodesics, demonstrating the atemporality mechanism. Section IV establishes linear stability and unitary quantum field propagation. Section V addresses topological aspects, causal structure, and thermodynamics. We conclude in Section VI with implications and future directions.
II The Signature-Changing BTZ Geometry and its Consistent Regularization
We develop the mathematical foundation for signature-changing black holes by constructing the Lorentzian-Euclidean BTZ metric and establishing a rigorous regularization scheme for its distributional curvature. This provides the consistent geometric framework necessary for analyzing singularity regularization.
II.1 The Lorentzian-Euclidean BTZ Metric
In standard BTZ coordinates $(t,r,\phi)$, the signature-changing metric takes the form:
$ds^{2}=-\varepsilon\left(-M+\frac{r^{2}}{\ell^{2}}\right)dt^{2}+\frac{dr^{2}}{-M+\frac{r^{2}}{\ell^{2}}}+r^{2}d\phi^{2},$ (1)
where the signature function $\varepsilon(r)$ is defined as:
$\varepsilon(r)=\text{sign}\left(-M+\frac{r^{2}}{\ell^{2}}\right)=2H\left(-M+\frac{r^{2}}{\ell^{2}}\right)-1,$ (2)
with $H(\cdot)$ representing the Heaviside step function normalized such that $H(0)=1/2$.
The metric exhibits distinctive features:
  • Exterior Region ($r>r_{h}$): For $r>\ell\sqrt{M}$, $\varepsilon=+1$, recovering standard Lorentzian BTZ.
  • Event Horizon ($r=r_{h}$): At $r=\ell\sqrt{M}$, the metric becomes degenerate: $\det(g_{\mu\nu})=0$.
  • Interior Region ($r<r_{h}$): For $r<\ell\sqrt{M}$, $\varepsilon=-1$, yielding Euclidean signature $(+,+,+)$.
To eliminate the coordinate singularity and facilitate physical interpretation, we introduce Painlevé-Gullstrand coordinates. For radial geodesics with normalization $g_{\mu\nu}u^{\mu}u^{\nu}=-\varepsilon$, the conserved quantities are:
$E$ $=\varepsilon^{2}\left(-M+\frac{r^{2}}{\ell^{2}}\right)\dot{t},$ (3)
$L$ $=r^{2}\dot{\phi}.$ (4)
For an observer falling from rest at infinity ($E=\varepsilon^{2}$, $L=0$), the radial equation simplifies to:
$\dot{r}=-\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}},$ (5)
with the coordinate relationship:
$\frac{dt}{dr}=\frac{\dot{t}}{\dot{r}}=-\frac{1}{\sqrt{\varepsilon}}\frac{1}{\sqrt{M-\frac{r^{2}}{\ell^{2}}}}\left(\frac{1}{-M+\frac{r^{2}}{\ell^{2}}}\right).$ (6)
Defining a new time coordinate $\mathscr{T}$ through $\sqrt{\varepsilon}\,\mathscr{T}(t,r):=\sqrt{\varepsilon}\,t+\Theta(r)$, where:
$\frac{d\Theta}{dr}=\frac{1}{-M+\frac{r^{2}}{\ell^{2}}}\sqrt{M-\frac{r^{2}}{\ell^{2}}},$ (7)
yields the Painlevé-Gullstrand form:
$ds^{2}=-\varepsilon d\mathscr{T}^{2}+\left(dr+\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}d\mathscr{T}\right)^{2}+r^{2}d\phi^{2}.$ (8)
This coordinate system remains finite at $r=r_{h}$ and provides the optimal framework for distributional analysis.
II.2 Distributional Curvature and Regularization Scheme
The presence of $\varepsilon(r)$ introduces distributional behavior in curvature tensors. Following Israel Israel (1966) and Barrabès-Israel Barrabes and Israel (1991), we define the jump discontinuity across $\Sigma$ as $F]:=F|_{+}-F|_{-}$. For our metric:
$[g_{\alpha\beta}]\neq 0,\quad[\partial_{\mu}g_{\alpha\beta}]\neq 0,\quad[\sigma_{AB}]=0.$ (9)
These discontinuities generate delta-function contributions to the Riemann tensor. The original regularization scheme Capozziello et al. (2024) employs a smooth approximation:
$\varepsilon(r)\rightarrow\frac{(r-r_{h})^{1/(2\kappa+1)}}{[(r-r_{h})^{2}+\rho]^{1/2(2\kappa+1)}},$ (10)
but encounters mathematical inconsistencies when handling terms like:
$\int dx\,\frac{x\varepsilon^{\prime\prime}(x)}{\varepsilon^{1/2}}=2\int dx\,\frac{x\delta^{\prime}(x)}{\varepsilon^{1/2}},$ (11)
where standard distributional identities fail due to singular behavior at $x=0$.
II.2.1 Modified Hadamard Regularization Scheme
We introduce a three-step modified Hadamard regularization:
Step 1: Controlled Regularization
$\varepsilon_{\rho}(x)=\frac{x^{1/(2\kappa+1)}}{(x^{2}+\rho)^{1/2(2\kappa+1)}},\quad x=r-r_{h},\quad\kappa\geq 2.$ (12)
Step 2: Consistent Distributional Treatment
  1. 1.
    Linear Terms: Employ Hadamard partie finie:
    $\int dx\,\frac{\delta(x)}{|x|^{n}}\equiv 0,\quad n>0.$ (13)
  2. 2.
    Nonlinear Terms: For sufficiently large $\kappa$:
    $\int dx\,f(x)\delta^{2}(x)=0\quad\text{when}\quad f(0)=0.$ (14)
  3. 3.
    Second Derivative Terms: Use identity:
    $\lim_{\rho\to 0^{+}}\int dx\,f(x)\varepsilon^{\prime\prime}_{\rho}(x)=\lim_{\rho\to 0^{+}}\int dx\,f^{\prime\prime}(x)\varepsilon_{\rho}(x).$ (15)
Step 3: Physical Validation - Ensure vanishing distributional contributions to $R_{\mu\nu}$ and agreement with standard BTZ geometry.
Figure 1: Regularized signature function $\varepsilon(r)$ for varying $\rho$ ($\kappa=1$ fixed). Larger $\rho$ (red) yields smoother transitions; smaller $\rho$ (green) approaches the discontinuous limit.
Figure 2: Regularized $\varepsilon(r)$ for varying $\kappa$ ($\rho=0.1r_{h}^{2}$ fixed). Lower $\kappa$ (red) gives sharper transitions; higher $\kappa$ (green) produces smoother transitions.
As shown in Figs. 1 and 2, our regularization provides a smooth transition between Lorentzian and Euclidean signatures, with parameters $\rho$ and $\kappa$ controlling the transition width and sharpness.
II.2.2 Application to BTZ Curvature
Applying our scheme yields the regularized Riemann tensor (detailed in Appendix A):
$R^{\alpha}_{\;\beta\mu\nu}=R^{\alpha}_{\;\beta\mu\nu\text{(reg)}}+\mathcal{O}(\rho^{(4\kappa-1)/2(2\kappa+1)}),$ (16)
with the regular part continuous except for finite jumps at $\Sigma$.
The resulting Ricci tensor satisfies:
$R_{\mu\nu}=0\quad\text{everywhere, including distributional sense},$ (17)
confirming no surface layers. The Weyl tensor contains no delta-functions, indicating no impulsive gravitational waves.
The regularized Kretschmann scalar remains finite:
$\mathcal{K}_{\text{reg}}=R_{\alpha\beta\mu\nu}R^{\alpha\beta\mu\nu}=\frac{12M^{2}}{\ell^{4}r^{4}}\left(r^{4}+\frac{2r^{2}\ell^{2}}{3}+\frac{\ell^{4}}{3}\right),$ (18)
attaining its maximum at the horizon: $\mathcal{K}_{\text{reg}}(r_{h})=\frac{16}{\ell^{4}}$.
This regularization scheme provides the mathematically consistent foundation for analyzing physical implications of signature change.
III Physical Analysis: Curvature and Singularity Avoidance
With the mathematical consistency established, we now examine the physical properties and implications of the signature-changing geometry, focusing on curvature structure, energy conditions, and the fundamental mechanism of singularity avoidance through atemporality.
III.1 Regularized Curvature and Energy Conditions
The complete set of regularized curvature tensors reveals the geometry’s physical structure:
III.1.1 Ricci Tensor and Scalar Curvature
$R_{\mathscr{T}\mathscr{T}}$ $=\frac{M}{\ell^{2}}\left(1-\frac{3r^{2}}{\ell^{2}}\right),$ (19)
$R_{rr}$ $=-\frac{M}{r^{2}\left(-M+\frac{r^{2}}{\ell^{2}}\right)}\left(1-\frac{r^{2}}{\ell^{2}}\right),$ (20)
$R_{\phi\phi}$ $=-\frac{Mr^{2}}{\ell^{2}},$ (21)
$R_{\mathscr{T}r}$ $=0,$ (22)
$R$ $=g^{\mu\nu}R_{\mu\nu}=-\frac{6M}{\ell^{2}}.$ (23)
The vanishing of distributional terms confirms no surface layers at $\Sigma$.
III.1.2 Weyl Tensor and Tidal Forces
$C^{\mathscr{T}}_{\;r\mathscr{T}r}$ $=\frac{M}{3\ell^{2}r}\left(1-\frac{3r^{2}}{\ell^{2}}\right),$ (24)
$C^{\phi}_{\;r\phi r}$ $=-\frac{M}{3r^{3}}\left(1-\frac{r^{2}}{\ell^{2}}\right),$ (25)
$C^{\mathscr{T}}_{\;\phi\mathscr{T}\phi}$ $=-\frac{Mr}{3\ell^{2}}\left(1-\frac{3r^{2}}{\ell^{2}}\right).$ (26)
No delta-function contributions exist, demonstrating absence of impulsive gravitational waves.
III.1.3 Energy Conditions and Effective Matter
Through $G_{\mu\nu}=8\pi T_{\mu\nu}^{\text{eff}}$, we obtain:
$T^{\text{eff}}_{\mathscr{T}\mathscr{T}}$ $=\frac{M}{8\pi\ell^{2}}\left(1-\frac{r^{2}}{\ell^{2}}\right),$ (27)
$T^{\text{eff}}_{rr}$ $=-\frac{M}{8\pi r^{2}\left(-M+\frac{r^{2}}{\ell^{2}}\right)}\left(1-\frac{r^{2}}{\ell^{2}}\right),$ (28)
$T^{\text{eff}}_{\phi\phi}$ $=-\frac{Mr^{2}}{8\pi\ell^{2}}.$ (29)
For null vectors $k^{\mu}=(1,\pm\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}},0)$:
$T^{\text{eff}}_{\mu\nu}k^{\mu}k^{\nu}=\frac{M}{8\pi\ell^{2}}\left(1-\frac{2r^{2}}{\ell^{2}}\right).$ (30)
The null energy condition is violated only deep inside the Euclidean region ($r<\frac{\ell}{\sqrt{2}}$), consistent with the exotic nature of signature-changing geometries.
III.2 Singularity Avoidance through Atemporality
The core achievement of signature change is the resolution of the central singularity via the dynamical mechanism of atemporality. This geometric regularization is further confirmed through the Kruskal-Szekeres extension developed in Appendix B, which provides a maximal analytic extension of the geometry and independently verifies the geodesic completeness results presented in this section.
III.2.1 Geodesic Motion and Proper Time Asymptotics
For radial geodesic motion with normalization $g_{\mu\nu}u^{\mu}u^{\nu}=-\varepsilon$:
$\dot{r}^{2}$ $=\varepsilon\left(M-\frac{r^{2}}{\ell^{2}}\right)-\varepsilon\left(1-\frac{E^{2}}{\varepsilon^{4}}\right),$ (31)
$\dot{t}$ $=\frac{E}{\varepsilon^{2}}\left(\frac{1}{-M+\frac{r^{2}}{\ell^{2}}}\right).$ (32)
The radial velocity reveals the atemporality mechanism:
$\dot{r}=-\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}.$ (33)
Critical features emerge:
  • As $r\to r_{h}^{+}$ ($\varepsilon\to 0^{+}$), $\dot{r}\to 0$
  • For $r<r_{h}$ ($\varepsilon=-1$), $\dot{r}$ becomes imaginary
  • Proper time to reach horizon diverges: $\sigma\to\infty$ as $r\to r_{h}$
Using parametrization $r(\eta)=r_{i}\cos^{2}(\eta/2)$, we derive:
$\frac{d\sigma}{d\eta}=r_{i}\sin\left(\eta/2\right)\cos^{2}\left(\eta/2\right)\sqrt{\frac{\varepsilon^{3}}{\varepsilon^{4}\sin^{2}\left(\eta/2\right)+E^{2}\left[\cos^{2}\left(\eta/2\right)-\varepsilon^{4}\right]}}.$ (34)
Refer to caption Refer to caption
(a) (b)
Figure 3: (a) Real proper time $\sigma(\eta)$ and (b) complete proper time for radial geodesic motion. Divergence at $\eta=\eta_{H}$ demonstrates infinite proper time to reach horizon.
As shown in Fig. 3(a), the proper time $\sigma(\eta)$ diverges as the observer approaches the horizon ($\eta\to\eta_{H}$), confirming that the horizon is reached only in the limit of infinite proper time. The complete behavior in Fig. 3(b) shows the imaginary continuation beyond the horizon, reflecting the emergence of atemporality.
Refer to caption Refer to caption
(a) (b)
Figure 4: (a) Coordinate time $t(\eta)$ in Lorentzian region. (b) Complete behavior showing Lorentzian (real) and Euclidean (imaginary) parts.
Fig. 4 illustrates the behavior of coordinate time. Panel (a) shows $t(\eta)$ diverging as the observer approaches the horizon in the Lorentzian region, while panel (b) displays the complete complex behavior, with the time coordinate becoming imaginary in the Euclidean interior.
Refer to caption Refer to caption
(a) (b)
Figure 5: (a) Comparison of $\sigma(\eta)$ and $t(\eta)$ in Lorentzian exterior. (b) Full comparison across both regions, showing imaginary proper time in Euclidean interior.
The comparison in Fig. 5 highlights the fundamental difference between proper time and coordinate time. While coordinate time diverges at the horizon, proper time approaches infinity, demonstrating that the horizon is physically unreachable in finite proper time.
III.2.2 Accelerated Observers and Universal Horizon Asymptotics
For radial acceleration with four-acceleration $a^{\lambda}$, the equations are:
$a^{t}$ $=\left(-M+\frac{r^{2}}{\ell^{2}}\right)^{-1}\frac{d}{d\sigma}\left[\left(-M+\frac{r^{2}}{\ell^{2}}\right)U^{t}\right],$ (35)
$a^{r}$ $=\frac{dU^{r}}{d\sigma}+\frac{M\varepsilon}{r^{2}}\left(-M+\frac{r^{2}}{\ell^{2}}\right)U^{t}U^{t}-\frac{M}{r^{2}\left(-M+\frac{r^{2}}{\ell^{2}}\right)}U^{r}U^{r}.$ (36)
The general solution shows universality:
$U^{t}$ $=\frac{\mathcal{F}}{-M+\frac{r^{2}}{\ell^{2}}},$ (37)
$U^{r}$ $=-\sqrt{\varepsilon}\sqrt{\mathcal{F}^{2}-\left(-M+\frac{r^{2}}{\ell^{2}}\right)},$ (38)
where $\mathcal{F}(\sigma)$ characterizes acceleration. For any finite $\mathcal{F}$ at $r=r_{h}$, $U^{r}\to 0$ as $r\to r_{h}$, becoming imaginary for $r<r_{h}$.
III.2.3 Physical Interpretation and Observational Implications
We interpret the imaginary radial velocity as dynamical emergence of atemporality:
  • Coordinate time $t$ becomes imaginary for $r<r_{h}$
  • Radial motion ceases to be parameterized by real proper time
  • Singularity at $r=0$ becomes causally disconnected
The bounded Kretschmann invariant (Eq. 18) combined with geodesic completeness demonstrates no physical singularities are accessible to observers, a conclusion further supported by the maximal analytic extension analysis.
Potential observational consequences include:
  1. 1.
    Particle Accumulation: Matter accumulation near $r=r_{h}$ could alter black hole silhouettes
  2. 2.
    Accretion Disk Structure: Modified interior dynamics may influence quasi-periodic oscillations
  3. 3.
    Quantum Gravity Probes: Elimination of central singularity suggests smoother black hole interiors
[…]
Figure 6: Penrose diagram for signature-changing BTZ geometry. Change surface $\Sigma$ separates Lorentzian exterior from Euclidean interior. Infalling geodesics asymptotically approach $\Sigma$ but cannot cross into Euclidean region.
The Penrose diagram in Fig. 6 illustrates the global causal structure of our geometry. The change surface $\Sigma$ acts as a one-way causal membrane, with infalling geodesics asymptotically approaching but never crossing into the Euclidean interior.
This comprehensive analysis establishes atemporality as a universal mechanism for singularity avoidance that operates for both geodesic and accelerated observers, while suggesting intriguing observational signatures for future investigation. The independent verification provided by the Kruskal-Szekeres extension in Appendix B strengthens the mathematical foundation of these results.
IV Robustness: Linear Stability and Quantum Field Propagation
A physically viable spacetime geometry must remain stable under perturbations and support consistent propagation of quantum fields. In this section, we perform a comprehensive stability analysis of the signature-changing BTZ geometry against gravitational perturbations and investigate the behavior of quantum scalar fields, serving as crucial tests for the physical robustness of our construction.
IV.1 Linear Stability Against Gravitational Perturbations
We consider small perturbations $h_{\mu\nu}$ around the background metric $g^{(0)}_{\mu\nu}$:
$g_{\mu\nu}=g^{(0)}_{\mu\nu}+h_{\mu\nu},\quad|h_{\mu\nu}|\ll 1.$ (39)
The linearized Einstein equations in vacuum, employing the transverse-traceless (TT) gauge $\nabla^{\mu}h_{\mu\nu}=0$, $h=0$, reduce to a wave equation for the perturbations:
[…]
For a radial perturbation of the form $h_{\mathscr{T}\mathscr{T}}=\psi(r)e^{-i\omega\mathscr{T}}$, we obtain a Schrödinger-like master equation:
$\frac{d^{2}\psi}{dr_{*}^{2}}+\left[\omega^{2}-V_{\text{eff}}(r)\right]\psi=0,$ (41)
where $dr_{*}=dr/(-M+r^{2}/\ell^{2})$ defines the tortoise coordinate.
The effective potential for gravitational perturbations in our geometry initially contains distributional terms:
$V_{\text{eff}}(r)=\varepsilon\left(M-\frac{r^{2}}{\ell^{2}}\right)\left[\frac{2M}{\ell^{2}r^{2}}+\frac{\varepsilon^{\prime}}{2r\varepsilon}\left(M-\frac{r^{2}}{\ell^{2}}\right)-\frac{\varepsilon^{\prime\prime}}{4\varepsilon}\left(M-\frac{r^{2}}{\ell^{2}}\right)\right].$ (42)
Applying our modified Hadamard regularization scheme (Section II.2) eliminates the singular contributions:
$\lim_{\rho\to 0^{+}}\int\frac{\varepsilon^{\prime}_{\rho}}{\varepsilon_{\rho}}f(r)dr=0,\quad\lim_{\rho\to 0^{+}}\int\frac{\varepsilon^{\prime\prime}_{\rho}}{\varepsilon_{\rho}}f(r)dr=0,$ (43)
yielding the regularized effective potential:
$V_{\text{eff}}^{\text{reg}}(r)=\varepsilon\left(M-\frac{r^{2}}{\ell^{2}}\right)\frac{2M}{\ell^{2}r^{2}}.$ (44)
The regularized potential $V_{\text{eff}}^{\text{reg}}(r)$ is non-negative in the Lorentzian domain ($\varepsilon=+1$, $r>r_{h}$), finite everywhere including at $r=r_{h}$, and vanishes as $r\to\infty$ and $r\to 0$. These properties ensure the positive-semidefiniteness of the operator $\frac{d^{2}}{dr_{*}^{2}}+V_{\text{eff}}^{\text{reg}}$ in the exterior, guaranteeing $\omega^{2}\geq 0$ and thus the absence of unstable modes ($\text{Im}(\omega)>0$) in the Lorentzian region.
To further corroborate stability, we employ the WKB approximation to estimate the quasi-normal mode frequencies. For gravitational perturbations, the first-order WKB formula yields:
$\omega^{2}=V_{0}-i\left(n+\frac{1}{2}\right)\sqrt{-2V_{0}^{\prime\prime}},$ (45)
where $V_{0}$ is the maximum of $V_{\text{eff}}^{\text{reg}}(r)$ in the Lorentzian region, occurring at $r=\sqrt{M}\ell$:
$V_{0}=\frac{2}{\ell^{2}},\quad V_{0}^{\prime\prime}=-\frac{8}{M\ell^{4}}.$ (46)
Substituting gives the gravitational quasi-normal mode spectrum:
$\omega_{n}^{\text{(grav)}}\approx\frac{\sqrt{2}}{\ell}-i\left(n+\frac{1}{2}\right)\frac{2}{\sqrt{M}\ell}+\mathcal{O}(1/n).$ (47)
We complement this analytical approach with a high-precision numerical computation using a Chebyshev spectral method (200 grid points, relative error $10^{-6}$). The numerical scheme carefully handles the transition at $r=r_{h}$ using our regularized potential, with boundary conditions requiring purely outgoing waves at spatial infinity and regularity at the change surface $\Sigma$ and the origin.
Refer to caption Refer to caption
(a) Quasi-normal mode spectrum (b) Regularized effective potential
Figure 7: (a) Numerical computation of quasi-normal modes for gravitational perturbations. All modes lie in the lower half-plane ($\text{Im}(\omega)<0$), indicating exponential decay and stability. Different markers represent different angular momentum numbers ($l=0,1,2$). (b) The regularized effective potential $V_{\text{eff}}^{\text{reg}}(r)$ showing its behavior across the change surface $\Sigma$ at $r=r_{h}$.
As shown in Fig. 7(a), all computed eigenvalues satisfy $\text{Im}(\omega)<0$, confirming the absence of unstable modes. The spectrum exhibits characteristic BTZ structure, with excellent agreement to WKB estimates for high overtones. Fig. 7(b) displays the regularized effective potential, which remains finite and well-behaved across the change surface.
Theorem 1.
The signature-changing BTZ geometry is linearly stable against gravitational perturbations. Comprehensive analytical and numerical analyses confirm that no exponentially growing modes exist. All perturbations remain bounded in time, and the quasi-normal mode spectrum exhibits characteristic exponential decay.
IV.2 Quantum Scalar Field Propagation
The behavior of quantum fields provides a fundamental test of causality, unitarity, and regularity. We investigate a massive scalar field $\phi$ with equation of motion:
[…]
In Painlevé-Gullstrand coordinates, the d’Alembertian operator is:
[…]
Employing separation of variables $\phi(\mathscr{T},r,\phi)=R(r)e^{-i\omega\mathscr{T}}e^{il\phi}$, the radial equation becomes:
$\frac{d}{dr}\left[r\left(-M+\frac{r^{2}}{\ell^{2}}\right)\frac{dR}{dr}\right]+\left[\frac{r\omega^{2}}{\varepsilon}-2\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}\omega r\frac{dR}{dr}-\frac{l^{2}}{r}-m^{2}r\right]R=0.$ (50)
The potentially problematic term is $\omega^{2}/\varepsilon$. Near the change surface $r=r_{h}+x$, with $\varepsilon_{\rho}(x)\sim x^{1/(2\kappa+1)}$, the radial equation becomes:
$\frac{d}{dx}\left[x\frac{dR}{dx}\right]+\frac{\omega^{2}r_{h}}{x^{1/(2\kappa+1)}}R+\text{regular terms}=0.$ (51)
IV.2.1 Singularity Structure via Frobenius Method
We analyze the singularity structure using the Frobenius method. Assuming a series solution $R(x)=x^{\alpha}\sum_{n=0}^{\infty}c_{n}x^{n}$, we obtain the indicial equation:
$\alpha(\alpha-1)+\alpha=\alpha^{2}=0\quad\Rightarrow\quad\alpha=0\quad(\text{double root}).$ (52)
The double root indicates a regular singular point. The general solution takes the form:
$R(x)=A\left[1+a_{1}x+a_{2}x^{2}+\cdots\right]+B\left[\ln x+b_{1}x+b_{2}x^{2}+\cdots\right].$ (53)
IV.2.2 Unitarity and Probability Conservation
Despite the logarithmic term in (53), unitarity—the conservation of probability—is preserved. The probability current for a complex scalar field is:
$j^{\mu}=-i\left(\phi^{*}\nabla^{\mu}\phi-\phi\nabla^{\mu}\phi^{*}\right).$ (54)
In our coordinates, the temporal component is:
$j^{\mathscr{T}}=-i\varepsilon\left(\phi^{*}\partial_{\mathscr{T}}\phi-\phi\partial_{\mathscr{T}}\phi^{*}\right)-i\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}\left(\phi^{*}\partial_{r}\phi-\phi\partial_{r}\phi^{*}\right).$ (55)
The crucial observation is that the factor $\varepsilon$ in the first term cancels the singular $\omega^{2}/\varepsilon$ behavior from the equation of motion. To verify finiteness explicitly, substitute the near-horizon expansion $\phi\sim(A+B\ln x)e^{-i\omega\mathscr{T}}$ into (55):
$j^{\mathscr{T}}$ $=-i\varepsilon_{\rho}(x)\left[(A^{*}+B^{*}\ln x)(-i\omega)(A+B\ln x)-(A+B\ln x)(i\omega)(A^{*}+B^{*}\ln x)\right]$
$\quad-i\sqrt{\varepsilon_{\rho}(x)}\sqrt{M-\frac{r_{h}^{2}}{\ell^{2}}}\left[(A^{*}+B^{*}\ln x)\left(\frac{B}{x}+\mathcal{O}(1)\right)-(A+B\ln x)\left(\frac{B^{*}}{x}+\mathcal{O}(1)\right)\right]$
$=2\omega\varepsilon_{\rho}(x)\left[|A|^{2}+2\text{Re}(A^{*}B)\ln x+|B|^{2}(\ln x)^{2}\right]$
$\quad-i\sqrt{\varepsilon_{\rho}(x)}\sqrt{M-\frac{r_{h}^{2}}{\ell^{2}}}\left[\frac{2i\,\text{Im}(A^{*}B)}{x}+\mathcal{O}(\ln x)\right].$ (56)
For $\kappa\geq 2$, $\varepsilon_{\rho}(x)\sim x^{1/(2\kappa+1)}$ vanishes faster than any negative power of $x$, ensuring both terms remain finite. Specifically:
  • First term: $\varepsilon_{\rho}(x)\times\text{polynomial in }\ln x\to 0$ as $x\to 0$
  • Second term: $\sqrt{\varepsilon_{\rho}(x)}/x\sim x^{1/(4\kappa+2)-1}=x^{-(4\kappa+1)/(4\kappa+2)}$ appears singular, but the numerator contains $\text{Im}(A^{*}B)$. By imposing the boundary condition $\text{Im}(A^{*}B)=0$ (which corresponds to choosing a real phase relationship between $A$ and $B$), this term vanishes identically.
The conservation law $\nabla_{\mu}j^{\mu}=0$ can be verified directly from the field equation. Integrating over a spacetime region and applying Gauss’s theorem shows that the total probability $P(\mathscr{T})=\int_{\Sigma_{\mathscr{T}}}\sqrt{h}\,n_{\mu}j^{\mu}\,d^{2}x$ is constant. Thus, despite the logarithmic term in the radial solution, probability is conserved and unitarity is maintained.
IV.2.3 Quasi-normal Mode Spectrum
For scalar field perturbations, following established BTZ results Cardoso and Lemos (2003); Horowitz and Hubeny (2000), the WKB-estimated quasi-normal mode spectrum is:
$\omega_{n}^{\text{(scalar)}}=\frac{\sqrt{M}}{\ell}\left[\pm l-i\left(n+\frac{1}{2}\right)\right]+\mathcal{O}(1/n).$ (57)
All modes exhibit $\text{Im}(\omega_{n})<0$, confirming exponential decay rather than growth, consistent with stability.
Theorem 2.
Quantum scalar field propagation in the signature-changing BTZ geometry is unitary and well-defined. The probability current remains finite and conserved across the change surface $\Sigma$, and all quasi-normal modes exhibit exponential decay ($\text{Im}(\omega)<0$), confirming stability.
IV.3 Physical Interpretation and Comparison
The numerical and analytical results provide strong evidence that the signature-changing modification introduces no pathologies:
  • Exterior Dynamics: The Lorentzian exterior matches the stable standard BTZ metric Birmingham and Sen (2001).
  • Interior Regularization: The Euclidean interior possesses bounded curvature and lacks the causal structure that typically drives instabilities.
  • Change Surface Behavior: Our regularization ensures smooth transition of perturbations across $\Sigma$.
  • Quantum Consistency: Scalar field propagation remains unitary and well-defined across the signature change, as demonstrated by probability current conservation (Section IV.2).
Comparison with the standard BTZ effective potential:
$V_{\text{eff}}^{\text{standard}}(r)=\left(-M+\frac{r^{2}}{\ell^{2}}\right)\frac{2M}{\ell^{2}r^{2}}$ (58)
reveals that our regularized potential (44) reduces to this form in the Lorentzian region ($\varepsilon=+1$), demonstrating continuity with the established stable case.
This comprehensive analysis establishes that the signature-changing BTZ geometry is robust against both classical and quantum perturbations, providing crucial evidence for its physical viability as a genuine vacuum solution that resolves the central singularity without introducing new instabilities.
V Interpretation and Consistency Checks
The consistent regularization of the signature-changing BTZ geometry necessitates a thorough examination of its physical implications and overall coherence. In this section, we address the fundamental nature of the $r=0$ region, the causal structure and interpretation of the Euclidean interior, and the thermodynamic properties of the black hole. We demonstrate that the central singularity is resolved into a milder topological boundary, the causal structure is well-defined and protected, and the external black hole thermodynamics remains unchanged, thereby reinforcing the physical viability of our construction.
V.1 Topological Nature of the $r=0$ Region and Geodesic Completeness
Our regularization scheme has demonstrated the finiteness of all curvature invariants, notably the Kretschmann scalar $\mathcal{K}_{\text{reg}}$ in Eq. (18), which remains bounded as $r\to 0$. This unequivocally eliminates the possibility of a curvature singularity at the origin.
To discern the true nature of the $r=0$ region, we examine the metric in the limit $r\ll\ell\sqrt{M}$. The Painlevé-Gullstrand metric (8) simplifies to:
$ds^{2}\approx-\varepsilon d\mathscr{T}^{2}+\left(dr+\sqrt{\varepsilon M}d\mathscr{T}\right)^{2}+r^{2}d\phi^{2}.$ (59)
Applying the transformation $T=\sqrt{\varepsilon}\mathscr{T}$ yields:
$ds^{2}\approx-dT^{2}+dr^{2}+2\sqrt{M}dTdr+r^{2}d\phi^{2}.$ (60)
The angular part, $r^{2}d\phi^{2}$, reveals no angular deficit:
$\Delta\phi=2\pi-\lim_{r\to 0}\frac{\text{Circumference}}{\text{Radius}}=2\pi-\lim_{r\to 0}\frac{2\pi r}{r}=0,$ (61)
ruling out a conventional conical singularity Deser et al. (1984).
However, consider a sequence of points $p_{n}$ with $r(p_{n})=1/n$. The proper distance between successive points is:
$d(p_{n},p_{n+1})=\int_{1/(n+1)}^{1/n}\frac{dr}{\sqrt{-M+\frac{r^{2}}{\ell^{2}}}}\approx\frac{1}{n(n+1)\sqrt{M}}.$ (62)
The series $\sum d(p_{n},p_{n+1})$ converges, indicating that $\{p_{n}\}$ is a Cauchy sequence. Yet, there exists no point $p$ in our manifold with $r(p)=0$ that serves as its limit.
Theorem 3.
The signature-changing BTZ spacetime is geodesically complete but metrically incomplete. The point $r=0$ constitutes a topological boundary singularity, a boundary point removed from the manifold, distinct from both curvature and conical singularities.
This represents a significant softening of the classical singularity. The pathological curvature divergence of the standard BTZ black hole, where $\mathcal{K}_{\text{BTZ}}\sim 1/r^{6}$, is replaced by a bounded curvature invariant and a manageable topological incompleteness.
V.2 Causal Structure and Interpretation of the Euclidean Interior
The change surface $\Sigma:r=r_{h}$ functions as a causal barrier with distinctive properties:
  • One-Way Causal Membrane: Future-directed causal curves cannot cross from the Lorentzian exterior ($r>r_{h}$) into the Euclidean interior ($r<r_{h}$).
  • Causal Disconnection: The Euclidean interior remains fundamentally inaccessible to external observers.
  • Chronology Horizon: The surface $\Sigma$ prevents causality violations in the observable Lorentzian domain, acting as a chronology protection boundary Hawking (1992).
This structure is manifest in the behavior of radial null geodesics:
$\frac{dr}{d\mathscr{T}}=-\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}.$ (63)
As $r\to r_{h}^{+}$ ($\varepsilon\to 0^{+}$), the velocity vanishes. For $r<r_{h}$ ($\varepsilon=-1$), the right-hand side becomes imaginary, signaling the breakdown of causal propagation and the emergence of atemporality.
We interpret this transition as a dynamical Wick rotation, physically realizing what is often a mathematical tool Sterman (1993). The metric undergoes a smooth signature change:
$ds^{2}=\begin{cases}-d\mathscr{T}^{2}+\cdots&(\varepsilon=+1,\ r>r_{h})\\ 0\cdot d\mathscr{T}^{2}+\cdots&(\varepsilon=0,\ r=r_{h})\\ +d\mathscr{T}_{E}^{2}+\cdots&(\varepsilon=-1,\ r<r_{h})\end{cases}$ (64)
where $\mathscr{T}_{E}=i\mathscr{T}$ emerges as the natural Euclidean time coordinate in the interior.
We posit that the Euclidean interior represents a frozen quantum gravitational phase. In this region:
  • Classical temporal dynamics ceases as time becomes a spacelike coordinate.
  • The geometry may encode entanglement between past and future states, providing a concrete geometric realization of the ER=EPR conjecture Maldacena and Susskind (2013).
  • It constitutes the gravitational analogue of the Hartle-Hawking ”no-boundary” initial condition Hartle and Hawking (1983b), transplanted to black hole physics.
The change surface $\Sigma$ thus demarcates the boundary between semiclassical Lorentzian evolution and a fundamentally quantum gravitational domain.
V.3 Black Hole Thermodynamics
A crucial test of physical consistency is whether the signature-changing interior alters the black hole’s thermodynamic properties as measured by an external observer.
V.3.1 Hawking Temperature and Surface Gravity
The Hawking temperature is determined by the surface gravity $\kappa_{g}$ at the event horizon:
$T_{H}=\frac{\kappa_{g}}{2\pi}.$ (65)
For our geometry, $\kappa_{g}$ is defined using the Killing vector $\xi^{\mu}=\delta^{\mu}_{\mathscr{T}}$:
$\kappa_{g}^{2}=-\frac{1}{2\varepsilon}(\nabla_{\mu}\xi_{\nu})(\nabla^{\mu}\xi^{\nu})\big|_{r=r_{h}}.$ (66)
A direct computation yields an expression containing a potentially ill-defined distributional term:
$\kappa_{g}=\left[\frac{\delta(1-r^{2}/(\ell^{2}M))}{\varepsilon(r)}\left(1-\frac{r^{2}}{\ell^{2}M}\right)+\frac{r}{\ell^{2}}\right]_{r=r_{h}}.$ (67)
Applying our modified Hadamard regularization scheme (cf. Section II.2):
$\int dr\,\frac{\delta(1-r^{2}/(\ell^{2}M))}{r}(r-r_{h})^{(2\kappa)/(2\kappa+1)}[(r-r_{h})^{2}+\rho]^{1/2(2\kappa+1)}=0,$ (68)
the distributional contribution vanishes, yielding the well-defined surface gravity:
$\kappa_{g}=\frac{r_{h}}{\ell^{2}}=\frac{\sqrt{M}}{\ell},$ (69)
and consequently, the Hawking temperature:
$T_{H}=\frac{\kappa_{g}}{2\pi}=\frac{\sqrt{M}}{2\pi\ell}.$ (70)
This result is identical to the standard BTZ black hole temperature Bañados et al. (1992).
V.3.2 Entropy and the First Law
The entropy follows from the first law of thermodynamics, $dM=T_{H}dS$. Substituting the temperature from Eq. (70):
$dS=\frac{dM}{T_{H}}=\frac{2\pi\ell}{\sqrt{M}}dM.$ (71)
Integration yields the Bekenstein-Hawking entropy:
$S=4\pi\ell\sqrt{M}=\frac{\pi r_{h}}{2G_{3}}=\frac{\mathcal{A}}{4G_{3}},$ (72)
where $\mathcal{A}=2\pi r_{h}$ is the horizon’s ”area” (circumference), and $G_{3}$ is the (2+1)-dimensional gravitational constant.
V.3.3 Free Energy and Phase Structure
The Helmholtz free energy is:
$F=M-T_{H}S=M-\frac{\sqrt{M}}{2\pi\ell}\cdot 4\pi\ell\sqrt{M}=-M.$ (73)
The negative free energy indicates the dominance of the black hole phase over pure thermal radiation in AdS space, mirroring the standard case. No new phase transitions are introduced by the signature change.
Theorem 4.
The thermodynamic properties of the signature-changing BTZ black hole—its Hawking temperature, Bekenstein-Hawking entropy, and free energy—are identical to those of the standard BTZ black hole. External thermodynamics is completely unaffected by the signature change in the interior.
This thermodynamic consistency underscores the robustness of our regularization scheme and aligns with the holographic nature of black hole thermodynamics, where external observables are insensitive to the detailed microstructure of the interior.
V.4 Summary of Physical Consistency
This unified analysis confirms the physical viability of the signature-changing BTZ geometry from multiple perspectives:
  • The central curvature singularity is resolved into a topologically incomplete boundary.
  • The causal structure is well-defined, with the change surface $\Sigma$ acting as a protective chronology horizon (Fig. 6).
  • Quantum field propagation and linear stability are maintained (Figs. 7).
  • The standard black hole thermodynamics is preserved for external observers.
These consistency checks collectively affirm that the paradigm of signature change, when implemented with mathematical rigor, provides a robust and physically plausible mechanism for classical singularity regularization.
VI Discussion and Conclusions
This work has established a mathematically rigorous foundation for signature-changing black holes by resolving critical inconsistencies in previous regularization schemes and demonstrating the viability of atemporality as a mechanism for singularity avoidance. Our comprehensive analysis of the $(2+1)$-dimensional BTZ geometry reveals that signature transition at the event horizon provides a natural mechanism for eliminating spacetime singularities without introducing unphysical distributional sources.
VI.1 Principal Achievements
Our investigation yields fundamental advances across mathematical consistency, geometric resolution, dynamical mechanisms, and physical viability:
VI.1.1 Mathematical Foundation of Signature Change
We identified and resolved a critical flaw in the original regularization scheme of Capozziello et al. Capozziello et al. (2024), where invalid manipulation of distributional terms involving $\varepsilon^{\prime\prime}(r)$ undermined the mathematical self-consistency of their vacuum solution. Our modified Hadamard regularization scheme provides:
  • A distributionally consistent treatment of second-order derivative terms
  • Rigorous elimination of spurious surface layers and impulsive gravitational waves
  • A robust framework for handling signature-changing geometries within established distribution theory
VI.1.2 Complete Geometric Resolution
The signature-changing BTZ geometry exhibits remarkable regularity and physical robustness, as demonstrated in Sections IIIV:
  • Vanishing distributional contributions to both Ricci and Weyl tensors with bounded curvature invariants throughout spacetime
  • Geodesic completeness established through Painlevé-Gullstrand and Kruskal-Szekeres extensions
  • Linear stability against gravitational perturbations and well-defined quantum field propagation (Fig. 7)
  • Resolution of the $r=0$ singularity into a topological boundary
  • Preservation of standard BTZ thermodynamics despite interior signature change
VI.1.3 Dynamical Atemporality Mechanism
We have transformed atemporality from a mathematical construct to a dynamical physical mechanism:
  • Radially infalling observers require infinite proper time to reach the horizon (Figs. 3, 5)
  • The transition to imaginary radial velocity provides a natural physical interpretation of Wick rotation
  • The Euclidean interior remains causally disconnected from the Lorentzian exterior (Fig. 6)
  • Universal operation for both geodesic and accelerated observers
VI.2 Theoretical Implications
VI.2.1 For Singularity Theorems
The bounded curvature and geodesic completeness demonstrated here challenge the inevitability conclusions of classical singularity theorems Hawking and Ellis (1973); Penrose (1969). Signature change provides a geometrically natural escape from the theorems’ assumptions through controlled energy condition violation in the Euclidean region, without requiring exotic matter sources.
VI.2.2 For Quantum Gravity
Our results suggest that some aspects of singularity regularization may be encoded in classical geometry rather than requiring full quantum gravity Rovelli (2004); Ashtekar and Lewandowski (2004). The bounded Kretschmann invariant indicates that curvature need not diverge at the Planck scale, potentially simplifying the interface between classical and quantum gravitational descriptions.
VI.2.3 For Black Hole Physics
The preservation of standard thermodynamic relations despite interior signature change reinforces the holographic nature of black hole thermodynamics, where external properties determine the thermodynamic description regardless of interior structure.
VI.3 Future Research Directions and Open Questions
Our work establishes a consistent framework for signature-changing black holes in $(2+1)$ dimensions, but several important extensions and open questions remain for future investigation.
VI.3.1 Mathematical and Dimensional Extensions
  • $(3+1)$-Dimensional Generalization: Applying our regularization scheme to Schwarzschild and Kerr geometries presents significant technical challenges due to more complex curvature structure and the presence of genuine curvature singularities. The extension would test whether signature change can similarly regularize four-dimensional singularities.
  • General Signature-Change Formalism: Developing a covariant formulation of signature change that goes beyond the BTZ ansatz, possibly through a smooth scalar field $\Phi$ with $g_{\mu\nu}\partial^{\mu}\Phi\partial^{\nu}\Phi$ changing sign, could provide a more fundamental description.
  • Distributional Geometry Rigor: A more general mathematical treatment of signature-changing metrics within Colombeau algebras or nonlinear distribution theories could further solidify the foundations.
VI.3.2 Physical and Phenomenological Investigations
  • Quantum Field Theory on Signature-Changing Backgrounds: Detailed study of quantum field propagation, vacuum polarization, particle production, and stress-energy tensor regularization in these geometries would clarify their semi-classical consistency.
  • Thermodynamic Refinements: Analysis of the third law, phase transitions between signature-changing and regular black holes, and potential modifications to the entropy-area relation due to the Euclidean interior.
  • Connection to Other Singularity Avoidance Mechanisms: Comparative study with regular black holes, noncommutative geometry, loop quantum gravity bounce mechanisms, and string theory fuzzballs could reveal unifying principles or distinctive features.
VI.3.3 Conceptual and Foundational Questions
  • Information Paradox Implications: If the interior is Euclidean and causally disconnected, how does this affect information retrieval? Does signature change naturally avoid firewalls or promote information conservation?
  • Causal Structure Interpretation: The precise interpretation of the change surface $\Sigma$ as a ”one-way causal membrane” requires deeper analysis within frameworks for non-globally hyperbolic spacetimes and quantum causality.
  • Relation to Quantum Gravity: While our approach demonstrates classical self-regularization, its connection to full quantum gravity—particularly whether the Euclidean interior represents a semiclassical approximation to a genuine quantum gravitational phase—remains open.
VI.3.4 Potential Observational Signatures
While challenging to detect, several potential empirical tests could emerge:
  • Gravitational Wave Modifications: Differences in quasi-normal mode spectra or ringdown behavior due to the modified interior structure, potentially detectable with next-generation gravitational wave observatories.
  • Black Hole Shadow Effects: Matter accumulation near the horizon (due to infinite proper time approach) could subtly alter shadow properties observable with the Event Horizon Telescope and future very-long-baseline interferometry.
  • Accretion Disk Dynamics: Modified interior geometry might influence quasi-periodic oscillations or other accretion phenomena in black hole X-ray binaries.
VI.3.5 Limitations and Caveats
It is important to acknowledge several limitations of our current work:
  • The $(2+1)$-dimensional setting, while mathematically tractable, lacks important physical features of realistic black holes (e.g., dynamical degrees of freedom, realistic matter coupling).
  • The mechanism prevents access to the singularity rather than eliminating it entirely; the $r=0$ region remains a topological boundary, though curvature remains finite.
  • While we demonstrate linear stability, nonlinear stability and long-term evolution remain to be investigated.
These directions highlight that signature change represents a promising but incomplete framework for singularity avoidance. Future work should address these extensions while maintaining the mathematical rigor established here, potentially bridging toward a more comprehensive understanding of spacetime singularities in classical and quantum gravity.
VI.4 Concluding Perspective
The signature-changing paradigm, when placed on firm mathematical footing through consistent regularization, offers a compelling alternative to quantum-gravitational singularity regularization. Our work demonstrates that classical general relativity possesses previously unrecognized capacity for self-regularization through fundamental geometric transformations.
The atemporality mechanism represents more than mere singularity avoidance—it provides a dynamical interpretation of the Euclideanization that has long been employed as a mathematical device in quantum gravity. By showing how signature change naturally emerges as a vacuum solution with bounded curvature and complete geodesics, we strengthen the case for considering geometric transformations as fundamental rather than phenomenological approaches to spacetime singularities.
As we continue to probe the extreme regimes of gravitational physics, signature-changing geometries stand as mathematically consistent and physically intriguing frameworks that may ultimately bridge our understanding of classical and quantum descriptions of spacetime.
Appendix A Regularization of the Riemann Tensor
In this appendix, we provide the detailed derivation of the regularized Riemann tensor components for the signature-changing BTZ metric. The complete calculation demonstrates the mathematical consistency of our modified Hadamard regularization scheme.
A.1 Christoffel Symbols and Their Distributional Behavior
We begin with the Painlevé-Gullstrand form of the BTZ metric:
$ds^{2}=-\varepsilon d\mathscr{T}^{2}+\left(dr+\sqrt{\varepsilon}\sqrt{M-\frac{r^{2}}{\ell^{2}}}d\mathscr{T}\right)^{2}+r^{2}d\phi^{2}.$ (74)
The non-vanishing Christoffel symbols contain distributional terms proportional to $\varepsilon^{\prime}(r)$ and $\varepsilon^{\prime\prime}(r)$:
$\Gamma^{\mathscr{T}}_{\mathscr{T}r}$ $=\frac{M}{r\left(-M+\frac{r^{2}}{\ell^{2}}\right)}+\frac{\varepsilon^{\prime}}{2\varepsilon},$ (75)
$\Gamma^{r}_{\mathscr{T}\mathscr{T}}$ $=\varepsilon\left(M-\frac{r^{2}}{\ell^{2}}\right)\left(\frac{M}{r\left(-M+\frac{r^{2}}{\ell^{2}}\right)}+\frac{\varepsilon^{\prime}}{2\varepsilon}\right),$ (76)
$\Gamma^{r}_{rr}$ $=-\frac{M}{r\left(-M+\frac{r^{2}}{\ell^{2}}\right)},$ (77)
$\Gamma^{r}_{\phi\phi}$ $=-r\left(-M+\frac{r^{2}}{\ell^{2}}\right),$ (78)
$\Gamma^{\phi}_{r\phi}$ $=\frac{1}{r}.$ (79)
The terms involving $\varepsilon^{\prime}$ and $\varepsilon^{\prime\prime}$ require careful regularization, as they generate Dirac-delta contributions in the curvature tensor.
A.2 Riemann Tensor Components Before Regularization
The Riemann tensor components calculated from the unregularized Christoffel symbols contain singular terms:
$R^{r}_{\;r\mathscr{T}r}$ $=\sqrt{\frac{M-\frac{r^{2}}{\ell^{2}}}{2}}\frac{r^{2}(2M-\frac{2r^{2}}{\ell^{2}})\varepsilon^{\prime 2}+2r\varepsilon\left[r(r-r_{h})\varepsilon^{\prime\prime}+3M\varepsilon^{\prime}\right]-8M\varepsilon^{2}}{r^{3}\varepsilon^{3/2}},$ (80)
$R^{r}_{\;\phi\phi r}$ $=\frac{M}{\ell^{2}},$ (81)
$R^{\mathscr{T}}_{\;r\mathscr{T}r}$ $=-\sqrt{\frac{2}{M-\frac{r^{2}}{\ell^{2}}}}R^{r}_{\;r\mathscr{T}r},$ (82)
$R^{\phi}_{\;r\phi r}$ $=\frac{M}{r^{3}\varepsilon}(r\varepsilon^{\prime}-\varepsilon).$ (83)
Similar singular behavior appears in the remaining components, all containing terms proportional to $\varepsilon^{\prime}$, $\varepsilon^{\prime 2}$, and $\varepsilon^{\prime\prime}$.
A.3 Step-by-Step Regularization Procedure
A.3.1 Step 1: Regularized Signature Function
We employ the regularized signature function:
$\varepsilon_{\rho}(x)=\frac{x^{1/(2\kappa+1)}}{(x^{2}+\rho)^{1/2(2\kappa+1)}},\quad x=r-r_{h},$ (84)
with $\kappa\geq 2$ to ensure the necessary differentiability.
A.3.2 Step 2: Treatment of Linear Delta-Function Terms
Consider the generic linear term:
$I_{1}=\int dx\,f(x)\frac{\delta(x)}{\varepsilon_{\rho}^{1/2}(x)}.$ (85)
Using the Hadamard partie finie prescription Hadamard (1923); Blanchet (2014):
$I_{1}=\lim_{\rho\to 0^{+}}\int dx\,f(x)\frac{\delta(x)(x^{2}+\rho)^{1/4(2\kappa+1)}}{x^{1/2(2\kappa+1)}}=0,$ (86)
since the integrand vanishes in the distributional sense due to the singular denominator.
A.3.3 Step 3: Treatment of Quadratic Delta-Function Terms
For terms quadratic in $\delta(x)$:
$I_{2}=\int dx\,f(x)\frac{\delta^{2}(x)}{\varepsilon_{\rho}^{3/2}(x)}.$ (87)
We observe that:
$\frac{\delta^{2}(x)}{\varepsilon_{\rho}^{3/2}(x)}=\delta^{2}(x)(x^{2}+\rho)^{3/4(2\kappa+1)}x^{(4\kappa-1)/2(2\kappa+1)}.$ (88)
For $\kappa\geq 2$, the coefficient vanishes at $x=0$, yielding $I_{2}=0$ in the distributional sense.
A.3.4 Step 4: Treatment of Second Derivative Terms
The most challenging terms involve $\varepsilon^{\prime\prime}_{\rho}(x)$. Consider:
$I_{3}=\int dx\,\frac{x\varepsilon^{\prime\prime}_{\rho}(x)}{\varepsilon_{\rho}^{1/2}(x)}.$ (89)
Instead of direct manipulation, we use the identity:
$I_{3}=\lim_{\rho\to 0^{+}}\int dx\,\frac{d^{2}}{dx^{2}}\left(\frac{x}{\varepsilon_{\rho}^{1/2}(x)}\right)\varepsilon_{\rho}(x).$ (90)
A detailed calculation shows:
$\frac{d^{2}}{dx^{2}}\left(\frac{x}{\varepsilon_{\rho}^{1/2}(x)}\right)=\mathcal{O}(x^{(4\kappa-3)/2(2\kappa+1)}),$ (91)
which, when combined with the behavior of $\varepsilon_{\rho}(x)$, gives $I_{3}\to 0$ as $\rho\to 0^{+}$ for $\kappa\geq 2$.
A.4 Final Regularized Riemann Tensor
After applying our regularization scheme to all components, we obtain the well-behaved Riemann tensor:
$R^{r}_{\;r\mathscr{T}r}$ $=-\frac{2M}{\ell^{2}r}\sqrt{\frac{2(M-\frac{r^{2}}{\ell^{2}})}{\varepsilon}},$ (92)
$R^{r}_{\;\phi\phi r}$ $=\frac{M}{\ell^{2}},$ (93)
$R^{\mathscr{T}}_{\;r\mathscr{T}r}$ $=\frac{2M}{\ell^{2}r\varepsilon},$ (94)
$R^{\phi}_{\;r\phi r}$ $=-\frac{M}{r^{3}},$ (95)
$R^{\phi}_{\;\mathscr{T}\phi r}$ $=\frac{M}{r^{2}}\sqrt{\frac{M-\frac{r^{2}}{\ell^{2}}}{2\varepsilon}},$ (96)
$R^{\mathscr{T}}_{\;\phi\mathscr{T}\phi}$ $=-\frac{M}{r}\left(M-\frac{r^{2}}{\ell^{2}}\right).$ (97)
The remaining components are either zero or related to these by symmetry.
A.5 Verification of Regularization Consistency
We verify three crucial properties of our regularized curvature:
  1. 1.
    Vacuum Solution: The regularized Ricci tensor satisfies $R_{\mu\nu}=0$ everywhere, including at $r=r_{h}$.
  2. 2.
    Finite Curvature Invariants: The Kretschmann invariant remains finite:
    $\mathcal{K}=R_{\alpha\beta\mu\nu}R^{\alpha\beta\mu\nu}=\frac{12M^{2}}{\ell^{4}r^{4}}\left(r^{4}+\frac{2r^{2}\ell^{2}}{3}+\frac{\ell^{4}}{3}\right).$ (98)
  3. 3.
    Distributional Consistency: All operations respect the principles of distribution theory, with no ill-defined products of distributions.
This appendix demonstrates that our modified Hadamard regularization scheme successfully resolves the mathematical inconsistencies in the original approach while preserving the essential physics of signature-changing black holes.
Appendix B Kruskal-Szekeres Extension for Signature-Changing BTZ Geometry
The Kruskal-Szekeres coordinates provide a maximal analytic extension of the BTZ geometry that removes the coordinate singularity at the event horizon while making the global causal structure manifest. In this appendix, we derive the Kruskal extension for the signature-changing BTZ metric and analyze its implications for the change surface $\Sigma$.
B.1 Coordinate Transformation
We begin with the Lorentzian-Euclidean BTZ metric in Schwarzschild-like coordinates:
$ds^{2}=-\varepsilon\left(-M+\frac{r^{2}}{\ell^{2}}\right)dt^{2}+\frac{dr^{2}}{-M+\frac{r^{2}}{\ell^{2}}}+r^{2}d\phi^{2},$ (99)
where $\varepsilon=\text{sign}(-M+r^{2}/\ell^{2})$.
Following the standard procedure for BTZ black holes Bañados et al. (1992), we introduce the tortoise coordinate:
$r^{*}=\frac{1}{\sqrt{\varepsilon}}\int\frac{dr}{-M+\frac{r^{2}}{\ell^{2}}}=\frac{\ell}{2\sqrt{\varepsilon M}}\ln\left|\frac{r-\ell\sqrt{M}}{r+\ell\sqrt{M}}\right|.$ (100)
We then define null coordinates:
$u=\sqrt{\varepsilon}t-r^{*},\quad v=\sqrt{\varepsilon}t+r^{*},$ (101)
in terms of which the metric becomes:
$ds^{2}=-\varepsilon\left(-M+\frac{r^{2}}{\ell^{2}}\right)du\,dv+r^{2}d\phi^{2}.$ (102)
The key step in the Kruskal extension is the introduction of new coordinates:
$U=-\frac{1}{\sqrt{\varepsilon}}e^{-\sqrt{\varepsilon}u/(2\ell\sqrt{M})},\quad V=\frac{1}{\sqrt{\varepsilon}}e^{\sqrt{\varepsilon}v/(2\ell\sqrt{M})}.$ (103)
In these coordinates, the metric takes the form:
$ds^{2}=-\frac{4\ell^{2}M}{\varepsilon}\frac{dU\,dV}{(1+\varepsilon UV)^{2}}+r^{2}d\phi^{2},$ (104)
where the radial coordinate $r$ is implicitly defined by:
$-\varepsilon UV=\frac{r-\ell\sqrt{M}}{r+\ell\sqrt{M}}e^{r/(\ell\sqrt{M})}.$ (105)
B.2 Analysis of the Change Surface
The change surface $\Sigma:r=\ell\sqrt{M}$ corresponds to the hypersurface where $\varepsilon=0$. In Kruskal coordinates, this occurs when:
$UV=0.$ (106)
The metric (104) reveals several important features:
  1. 1.
    Regularity at Horizon: The metric components remain finite at $UV=0$, confirming that the coordinate singularity has been removed.
  2. 2.
    Signature Change: The factor $1/\varepsilon$ in the metric indicates that the geometry becomes degenerate at $\Sigma$, consistent with our findings in Painlevé-Gullstrand coordinates.
  3. 3.
    Maximal Extension: The coordinates $(U,V)$ cover the entire maximally extended spacetime, with four regions:
    • Region I: $U<0$, $V>0$ (Lorentzian exterior)
    • Region II: $U>0$, $V>0$ (Lorentzian interior)
    • Region III: $U<0$, $V<0$ (Euclidean exterior)
    • Region IV: $U>0$, $V<0$ (Euclidean interior)
B.3 Comparison with Standard BTZ Extension
The Kruskal extension for the signature-changing BTZ geometry differs from the standard BTZ case in several crucial aspects:
  • Signature Dependence: The coordinate transformation explicitly depends on $\sqrt{\varepsilon}$, which becomes imaginary in the Euclidean regions.
  • Causal Structure: The usual bifurcate horizon structure is modified. Instead of two intersecting null surfaces, we have a single change surface $\Sigma$ that separates Lorentzian and Euclidean regions.
  • Analytic Continuation: The metric (104) suggests that the Euclidean and Lorentzian regions are related by analytic continuation in the time coordinate, but with the important distinction that the signature change occurs dynamically rather than as a mathematical artifact.
B.4 Geodesic Completeness in Kruskal Coordinates
The behavior of geodesics in Kruskal coordinates provides additional insight into the atemporality mechanism:
Theorem 5.
All causal geodesics in the signature-changing BTZ geometry are complete in Kruskal coordinates, with no geodesic reaching the $r=0$ singularity in finite affine parameter.
Proof.
Consider a radial null geodesic with tangent vector $k^{\mu}$. In Kruskal coordinates, the geodesic equation reduces to:
$\frac{d^{2}U}{d\lambda^{2}}=0,\quad\frac{d^{2}V}{d\lambda^{2}}=0,$ (107)
where $\lambda$ is an affine parameter. The solutions are linear functions:
$U(\lambda)=U_{0}+A\lambda,\quad V(\lambda)=V_{0}+B\lambda.$ (108)
From equation (105), the $r=0$ singularity corresponds to:
$-\varepsilon UV=-1.$ (109)
For any finite values of the constants $A$, $B$, $U_{0}$, $V_{0}$, the product $UV$ remains finite for all finite $\lambda$. Therefore, no geodesic reaches $r=0$ in finite affine parameter, demonstrating geodesic completeness. ∎
B.5 Global Structure and Penrose Diagram
The Penrose diagram derived from the Kruskal extension (Fig. 6) exhibits a novel structure:
  • The change surface $\Sigma$ appears as a null line separating Lorentzian and Euclidean regions.
  • The diagram is not square, reflecting the different causal structures in the two regions.
  • There are no Cauchy horizons, and the initial value problem remains well-posed when restricted to the Lorentzian domain.
B.6 Implications for Regularization
The Kruskal extension provides independent confirmation of our regularization results:
  1. 1.
    Curvature Regularity: The Riemann tensor components computed in Kruskal coordinates remain finite everywhere except at $r=0$, consistent with our Painlevé-Gullstrand analysis.
  2. 2.
    Distributional Terms: The coordinate transformation to Kruskal coordinates is smooth everywhere except at $\Sigma$, where it becomes degenerate. This degeneracy is the source of the distributional terms we regularized in Section II.2.
  3. 3.
    Physical Interpretation: The maximal extension suggests that the signature-changing BTZ geometry represents a physically reasonable vacuum solution, with the change surface acting as a natural boundary between causally disconnected regions.
This Kruskal analysis complements our Painlevé-Gullstrand approach and provides additional evidence for the mathematical consistency and physical viability of signature-changing black hole geometries.